### v1

Abstract
We analyze blocks proposed for inclusion in the Ethereum blockchain during 8 minutes on December 3rd, 2024. Our dataset comprises 38 winning blocks, 15,097 proposed blocks, 10,793 unique transactions, and 2,380,014 transaction-block pairings. We find that exclusive transactions—transactions present only in blocks proposed by a single builder — account for 85% of the fees paid by all transactions included in winning blocks. We also find that a surprisingly large number of user transactions are delayed: although proposed during a bidding cycle, they are not included in the corresponding winning block. Many such delayed transactions are exclusive to a losing builder. We also identify two arbitrage bots trading between decentralized (DEX) and centralized exchanges (CEX). By examining their bidding dynamics, we estimate that the implied price at which these bots trade USDC/WETH and USDT/WETH on CEXes is between 3.4 and 4.2 basis points better than the contemporaneous price reported on Binance.
Keywords:
Ethereum, Proposer-Builder Separation (PBS), Order Flow.
1 Introduction
inlineinlinetodo: inlinebefore submitting, we should make sure that we are consistent with the naming of builders and relaysinlineinlinetodo: inlinechange ”block value” to ”block’s potential revenues” (or somethign else)
Blockchains are often described as immutable public ledgers because transactions included in the canonical chain cannot be altered. This static characterization, however, conceals the probabilistic and competitive process that determines which transactions become immutable. In Ethereum, for each block added to the canonical chain, up to 5,000 candidate blocks are constructed, propagated, but ultimately discarded along with their transactions. Each of these discarded blocks represents a counterfactual, unrealized history. In one candidate block, a user might secure a highly profitable trade; in another, the same user could suffer a loss due to an attack. When such differences arise, the choice of winning block has significant consequences for users. Conversely, if blocks differ only in a handful of arbitrage transactions, the outcome may matter little to regular users. Understanding how proposed blocks differ and the resulting implications for users is essential to understanding a system that facilitates roughly USD 5 billion in daily financial exchanges.
In this paper, we perform the first analysis of transactions included in non-winning Ethereum blocks. We collect 15,097 non-winning blocks submitted between Ethereum blocks 21,322,622 to 21,322,660 (38 winning blocks), approximately from 2:37:35 PM to 2:45:25 PM UTC on December 3, 2024. Our dataset contains 10,793 individual transactions (identified by their hash) and 2.380.014 individual transactions/blocks combinations, which will be our main unit of analysis. During the approximately eight-minute we consider, the price of ETH against USDC on Binance showed notable volatility. It began at 3,517.45 (2:37:35 PM UTC), rose by 1.06% to 3,554.92 (2:42:12 PM), dipped by 0.37% to 3,541.68 (2:43:13 PM), and then increased again by 0.60% to close at 3,562.79 (2:45:25 PM) (see Figure 1).
Before presenting our main results, we review the process by which blocks are submitted and selected for inclusion (readers familiar with it may skip this section) and describe our dataset in greater detail, including its limitations. After presenting our findings, we discuss their relevance to the broader debate about Ethereum’s current design and their relation to the existing academic literature.
Figure 1: ETH/USDC price on Binance during our study period
1.1 Background: the journey of an Ethereum transaction
On Ethereum, the right to add a new block to the canonical chain is assigned to a proposer (or validator), randomly selected from addresses that have locked (or staked) a sufficient amount of ETH, Ethereum’s native token. The proposer can assemble the block using transactions from the public mempool, a repository of pending transactions submitted by Ethereum users. However, in over 90% of cases, the proposer delegates block creation to specialized entities called builders, who gather both user transactions and those submitted by searchers. Searchers are automated bots that generate transactions in response to market conditions (for example, price fluctuations) or in response to other pending transactions in the mempool. For builders, selecting which transactions to include and their order is a complex optimization problem: block space is limited, not all transactions can be included, and some may conflict with others.
Users and searchers interact with builders in different ways. Some users submit their transactions to the Ethereum public mempool, where builders observe them. However, this exposes them to potential attacks. [^1] For example, if a user swaps assets on a blockchain-based financial market (an Automated Market Maker, or AMM), a malicious searcher can execute a “sandwich” attack: it front-runs the user’s swap by making the same trade first, then back-runs it with the opposite swap. This allows the attacker to buy low and sell high, forcing the victim to trade at worse prices. To mitigate this risk, users may instead use private mempools, services that connect them directly with multiple builders. Searchers always submit their transactions privately, either through known private mempools or via direct communication with one or more builders. Every transaction includes either a fee or a payment to the “Fee Recipient” address, a block-level variable set by the builder (typically an Ethereum address they control). As a result, builders earn fees from users and searchers whenever they include their transactions in a block that is added to the blockchain.
The relation between builders and proposers is facilitated by relays. Their primary role is to prevent proposers from altering the submitted blocks, for example, by changing the “Fee Recipient” variable. [^2] Technically, this is achieved by requiring proposers to sign an empty block, which is then completed by the relay. For more details, see https://docs.flashbots.net/flashbots-mev-boost/relay. Once a builder creates a block, it submits it to a relay along with a proposed payment for the proposer. The relay then forwards the highest-paying block and its corresponding payment to the proposer upon request. This process resembles an ascending price auction because the relay continuously broadcasts the value of the highest-paying bid—though this broadcast is observed with a delay due to network latency. An auction cycle lasts approximately 12 seconds, after which the winning block is selected, and a new auction begins. During each auction cycle, builders continuously resubmit blocks, modifying the transactions included and bid amounts. Latency plays a crucial role, as proposers are geographically dispersed, and arbitrage transactions often depend on rapid price movements in traditional “off-chain” financial markets. To account for geographic dispersion and fast-moving arbitrage opportunities, multiple relays operate worldwide, with builders submitting to several at once. As a result, thousands of blocks are created and discarded for every block added to the Ethereum blockchain.
1.2 Our dataset
Our primary dataset includes all blocks (winning and non-winning) submitted via a relay called “Agnostic relay” during the study period. [^3] Founded in 2022, Agnostic relay held 20% of the relay market in 2023, declining to around 5% during our study period. We complement our primary dataset with several publicly available data sources:
  • Winning blocks for our study period, including those not supplied by Agnostic relay. Additionally, we track 10 winning blocks beyond our study period to determine whether transactions in our primary dataset were eventually included in the blockchain.
  • Hashes of blocks submitted to other relays (Flashbots, BloXroute, Manifold, Eden, Ultra Sound, SecureRpc, and Aestus), along with the name of the builders submitting them and their bids. Here, a block’s hash serves as its unique identifier, as two blocks that differ in a single transaction or their bids have different hashes. We can, therefore, track whether a block in our primary dataset was also submitted to other relays and its associated bid. In particular, we find that the 15,097 submitted blocks included in our primary dataset constitute 28.3% of all blocks submitted via major relays during the study period.
  • Data on transactions submitted via the public mempool and two known private mempools (MEV Blocker and Flashbots Protect), obtained from https://mempool.guru/ and https://Dune.com.
  • Second-by-second price data from Binance.
We also simulate the non-winning blocks in our dataset to assess how the execution of a transaction — particularly a swap on a blockchain-based financial market — varies depending on the block in which it is included. We simulate the block using an Ethereum Archive Node running the Erigon client, forking at the appropriate block heights, and replaying the blocks.
Our primary dataset has two main limitations. First, we observe no submitted blocks for 11 bidding cycles because the proposers assigned to those slots were not connected to the Agnostic relay, which therefore did not collect block submissions. Second, it contains proportionally few blocks from Beaverbuild, a major block builder during the study period: while Beaverbuild accounted for 40.2% of all blocks submitted to major relays, it contributed only 9.4% of the blocks submitted to Agnostic (and thus included in our dataset). We later discuss how latency may explain Beaverbuild’s limited presence on Agnostic. With these limitations in mind, we can now discuss our main results.
1.3 Summary of the main results.
The first observation is that the set of proposed transactions differs greatly from the subset of these transactions that appear in winning blocks. The vast majority (87.7%) of the transactions in our primary dataset that appear in winning blocks are users’ transactions, and only a few are searchers’ transactions. [^4] We classify an address as a searcher if it trades on decentralized exchanges (DEXes) using a smart contract whose source code is not disclosed on https://Etherscan.com. In other words, part of the searcher’s logic is embedded in a smart contract but hidden by making only its bytecode available (the bytecode is the compiled, machine-readable version of the smart contract). Swapping on a DEX is necessary for all common types of “searching,” such as sandwich attacks, arbitrage, backrunning, and triggering liquidations on lending protocols. Also, only few (14.4%) of users’ transactions are swaps. Instead, among the set of proposed transactions, about half are from users and the rest from searchers. Also, about half of the proposed transactions are swaps, which are highly concentrated among a few searchers: the top four addresses alone account for 3,653 swap transactions, or approximately 70% of the total swaps.
A notable feature of our dataset is that 2,981 transactions — 35% of all unique transactions — originate from just two addresses that almost exclusively perform swaps. We identify these addresses as searchers integrated with the two main builders in our dataset, Rsync and Titan, and refer to them as “Rsync-bot” and “Titan-bot,” respectively. Titan-bot sends transactions exclusively to Titan, while Rsync-bot sends transactions mainly to Rsync but also to Titan. By examining transaction logs, we find that all Rsync-bot transactions sent to Titan were also sent to Rsync. Notably, in every instance where the Rsync-bot submits the same transaction to both builders, the fee offered to Titan is always lower than the one offered to Rsync, with an average difference of 18%. We interpret this gap as the bot’s target profit margin when interacting with builders that are not integrated with it.
Our dataset allows us to distinguish between private transactions shared among multiple builders and exclusive transactions — those available to a single builder. Among the 5,576 transactions in our primary dataset that were included in a winning block, Mempoolguru classifies 4,124 as public and 1,452 as non-public. For the non-public subset, we check whether each transaction appeared in blocks proposed by multiple builders (private) or only in blocks from a single builder (exclusive) during the bidding cycle that led to their inclusion on chain. We find that 45% of non-public transactions included in winning blocks are exclusive. These transactions are responsible for 85% of the winning blocks’ revenues (i.e., the sum of the payments to the builder), while private but not exclusive transactions constitute 10% of revenues.
Interestingly, we find a number of transactions that change status over time: from exclusive to private and even from exclusive to public. Specifically, we identify 19 transactions that are exclusive to a single builder during the initial part of the auction cycle and later appear in blocks from other builders, thereby transitioning from exclusive to private within the same auction cycle. In the majority of cases, these transactions are submitted via Flashbots Protect and are initially exclusive to the Flashbots builder before being shared with others. We also observe 12 users’ transactions that appear exclusive to a single builder for one or two bidding cycles, and later appear in the public mempool when the builder does not win. We speculate that a private mempool operator shares each transaction exclusively with a single builder, falling back to the public mempool if that builder does not win. [^5] The presence of transactions that change status is noteworthy because four of the five major private mempool operators explicitly state that they share transactions with multiple builders (those are MEV Blocker, Flashbots Protect, Blink, and Merkel; for an analysis of the difference between these operators see Janicot and Vinyas, 2025). In contrast, the fifth major private mempool operator (Metamask Smart Transactions) does not publicly document its sharing policies.
One of our main findings is that approximately 20% of users’ transactions are delayed: they are proposed during a bidding cycle before the winning block is chosen, but they are not included in the corresponding winning block. Strikingly, 30% of these delayed transactions initially appear as exclusive. This is difficult to rationalize, as users do not typically trade based on superior information, and hence gain no advantage from sending transactions exclusively to a single builder rather than sharing them privately with multiple builders or broadcasting them publicly. Even more surprisingly, as already discussed, some of these exclusive transactions later appear in the public mempool. These users’ transactions are therefore delayed and also exposed to attacks.
We also examine what determines whether a swap executes in one block but not another, and the price at which it executes. Regression results show that a swap in the same direction as Rsync-bot or Titan-bot is approximately 18% less likely to execute when included in a block built by Rsync or Titan that also includes a transaction from their respective bot. In contrast, a swap in the opposite direction is 1% more likely to execute under the same conditions. A similar, albeit smaller, pattern emerges for prices: swaps in the opposite direction as the two bots trade at better prices when included in blocks by either Titan or Rsync that also include transactions of their respective bots, while the opposite is true for swaps in the same direction as the bots.
Finally, we study the competition between Rsync-bot and Titan-bot. Because each bot is integrated with its respective builder, its fee for block inclusion measures its risk-adjusted expected profit. By observing each bot’s swap volume, the on-chain execution price, and its inclusion payment, we infer an implied centralized exchange (CEX) price for the off-chain leg of the trade. This allows us to compare the bots’ implied CEX prices to the contemporaneous Binance price and examine how this difference evolves during the bidding cycle. Focusing on instances where both bots compete to rebalance either a USDT/WETH pool or a USDC/WETH pool, we find that the implied CEX price for a marginal (zero-volume) trade at second 12 of the bidding cycle is approximately 2.3 basis points better than Binance’s price for Rsync-bot and 1.5 basis points better for Titan-bot. For a 10 ETH trade, the corresponding improvements over Binance are 2 basis points for Rsync-bot and 1.2 for Titan-bot. For other token pairs, for which we have much fewer observations than for USDT/WETH and USDC/WETH, the evidence is ambiguous and we find no indication that the bots trade at a better or worse price than the one reported by Binance.
Finally, we analyze the competition between Rsync-bot and Titan-bot. Since each bot is integrated with its respective builder, its payment for block inclusion equals its risk-adjusted expected profit. By observing each bot’s swap volume, on-chain execution price, and payment for inclusion, we infer an implied centralized exchange (CEX) price — i.e., the price of the off-chain leg of the trade. This inferred CEX price can then be compared to the contemporaneous Binance price during the bidding cycle. Focusing on cases where both bots compete to rebalance either a USDT/WETH or a USDC/WETH pool, we find that the implied CEX price is approximately 4.2 basis points better than Binance’s price for Rsync-bot, and 3.4 basis points better for Titan-bot.
1.4 Contribution
Ethereum’s process of block creation, submission, and selection— known as Proposer-Builder Separation (PBS) — has been the subject of considerable debate and criticism. [^6] For a summary of this debade, with a particular emphasis on the economic aspects, see John et al. (2025). Our results contribute directly to these discussions.
A primary concern is the extreme concentration in the builders’ market: the top two builders (Titan and Beaverbuild) currently produce between 80% and 90% of blocks added to the Ethereum blockchain. These entities wield substantial influence over transaction inclusion, posing a threat to the blockchain’s foundational goals of decentralization and permissionlessness. This concentration is often attributed to economies of scale from exclusive transactions: the more often a builder wins, the more likely it is to attract transactions not shared with competitors, and the more this builder wins. Until now, this hypothesis has remained untested, as on-chain data alone does not allow distinguishing between transactions shared privately with multiple builders and truly exclusive ones. Our analysis fills this gap by showing that 85% of the revenues from winning blocks in our dataset, measured by total payments to builders, is generated by exclusive transactions.
A second major criticism of PBS is that it enables searchers to efficiently extract value from users. Quantitatively, the main concern is arbitrageurs profiting at the expense of on-chain liquidity providers by exploiting price discrepancies between decentralized exchanges (DEXes) and centralized exchanges (CEXes). The resulting loss to liquidity providers is referred to as Loss-vs-Rebalancing (LVR), a concept introduced by Milionis et al. (2022). Subsequent studies have attempted to identify arbitrageurs (Heimbach et al., 2024) and quantify LVR (Canidio and Fritsch, 2023; Fritsch and Canidio, 2024). These analyses, however, rely on strong assumptions. This is because on-chain data alone reveals only that a swap occurred, not whether it was part of an arbitrage strategy. Furthermore, the prices at which arbitrageurs trade on centralized venues are usually unobservable. Our dataset overcomes these limitations in two ways. First, we can identify arbitrage bots by observing competition between them for the same arbitrage opportunities. [^7] In this respect, our approach resembles that of Aquilina et al. (2022), who identify high-frequency trading races in traditional financial markets by analyzing failed transactions. Second, by focusing on builder-integrated searchers engaging in CEX-DEX arbitrage, we can infer their expected profits and thus back out the effective price at which they trade on the centralized leg.
A final criticism is that a large section of the PBS supply chain is opaque and may not operate in the users’ interest. Our paper illuminates some of its aspects and shows that, indeed, there are reasons for concern regarding the handling of users’ transactions.
1.5 Additional relevant literature
The dynamics of builder competition and block inclusion on the Ethereum blockchain have been studied extensively. However, to the best of our knowledge, we are the first to systematically collect data on transactions contained in non-winning blocks. Yang et al. (2024) is the only other paper that analyzes the content of non-winning blocks. They examine competition among builders by comparing builders’ bids to the blocks’ revenues. Several other papers explore different aspects of builder competition. For example, Wu et al. (2024a) and Wu et al. (2024b) study builders’ bidding strategies, both theoretically and through simulation. Titan and Frontier Research (2023) and Öz et al. (2024) show that builder profits depend on how they source transactions, using a classification approach similar to ours. Bahrani et al. (2024) analyze builder competition theoretically and derive conditions under which the builder and proposer markets become concentrated in equilibrium. While we also study builder competition, our focus is on block-building strategies, for instance, whether they are vertically integrated with searchers and the behavior of these searchers.
Similarly to our analysis of the competition between Rsync-bot and Titan-bot, Capponi et al. (2024) also studies searchers competing to exploit CEX-DEX arbitrage opportunities. However, they consider an earlier period in which searchers competed by submitting transactions to the public mempool and engaging in priority fee auctions — i.e., because all transactions are included in the block and ordered by priority fee, the winning searcher is the one paying the highest fee. The current market structure and nature of the competition between searchers are very different, with searchers submitting directly to builders, builders selecting which searchers’ transactions to include, and frequent integration between searchers and builders. [^8] Also notable is Capponi et al. (2023), in which the same authors develop a theory of searchers’ competition when transactions are sent privately to builders.
Also related are Öz et al. (2023) and Schwarz-Schilling et al. (2023), who study how proposers decide when to request the winning block. They argue that this choice is partially strategic, as builders’ bids tend to increase during the auction cycle. Delaying the selection of the winning block can, therefore, lead to higher payments to the proposer. Wahrstätter et al. (2023) and Heimbach et al. (2023) provide an empirical analysis of Ethereum block production from September 2022 to May 2023, a period during which proposer-builder separation (PBS) became the dominant mechanism for block construction on Ethereum.
Finally, Pai and Resnick (2024) develops a theory of integration between builders and searchers, based on the idea that an integrated searcher’s payment to a builder reflects the full value of including the transaction. In contrast, non-integrated searchers pay less than the full value to retain a profit if their transaction is included. This has implications for builder competition, as it suggests that builders integrated with searchers enjoy a competitive advantage over non-integrated builders. These insights are relevant to our analysis. In particular, we observe integrated searchers submitting the same arbitrage transaction both to the builder with whom they are integrated and to others with whom they are not. In such cases, the payment offered to the integrated builder is consistently higher than that offered to non-integrated builders. We interpret this higher payment as reflecting the true value of transaction inclusion.
The remainder of the paper is organized as follows. The next section describes how we constructed our dataset and provides additional summary statistics. The following section provides an in-depth analysis of the auction cycle leading to the addition of Block 21322649 to the Ethereum blockchain. We then extend our analysis to all auction cycles in our dataset. The last section concludes.
2 Dataset Description and summary statistics
2.1 Blocks
Ethereum subdivides time into 12-second slots, during which one block at most can be added to the blockchain. For our purposes, each slot corresponds to a bidding cycle. Slot numbers differ from block numbers; for example, our dataset covers all blocks submitted via the Agnostic relay from slot 10,534,387 to 10,534,425, corresponding to winning blocks 21,322,622 to 21,322,660. For ease of exposition, in what follows, we use winning block numbers whenever possible. Also relevant is that a relay will not hold an auction for a given slot if the proposer assigned to that slot is not registered with the relay. This situation appears in our data: out of the 39 slots we consider, we observe no bids for 11 bidding cycles because the proposers were not connected with the Agnostic relay. [^9] Specifically, we observe no bids for the auction cycles that led to the inclusion of blocks 21,322,624; 21,322,625; 21,322,627; 21,322,629; 21,322,633; 21,322,634; 21,322,636; 21,322,644; 21,322,646; 21,322,647; 21,322,651.
Each block in our dataset has two timestamps: received at (the time the relay received the block from the builder) and made available at (the time the relay made the block available to the proposer). The difference between these timestamps arises because relays simulate blocks to verify their validity before making them available to the proposer. [^10] Some blocks are treated “optimistically” and made available to the proposer immediately, with simulation performed in the background. If the simulation later fails, the builder’s next submission will not be treated optimistically. In our dataset, such “optimistic” blocks constitute only 3.8% of the sample. The simulation delay is non-negligible: the median time between the two timestamps is 0.76 seconds, with the 75th percentile at 1.5 seconds. In what follows, we focus on the received at timestamp, as it more accurately reflects the key metrics we study, such as when a builder submits a new block or when a transaction first appears in a block.
As previously noted, our primary dataset includes 15,097 blocks submitted across 28 bidding cycles. The number of blocks per slot ranges from a minimum of 220 to a maximum of 951, with a mean of 539.18 and a median of 509. We are able to match 14,043 of these blocks to 23 known builders. The most active builders in our dataset are Titan Builder (7,024 blocks, 46.5%), Rsync Builder (2,259 blocks, 15.6%), Flashbots (1,936 blocks, 12.8%), and Beaverbuild (1,418 blocks, 9.4%). There is a substantial overlap between blocks submitted to Agnostic and those submitted to other relays: 12,895 blocks in our primary dataset were also shared with at least one other relay.
However, when examining the proportion of blocks submitted by the top builders across all relays, we note an imbalance in our primary dataset. Across all relays, Titan Builder, submitted 13,679 blocks (25.6%), Rsync Builder 7,936 blocks (14.9%), Flashbots 2,626 blocks (4.9%), and Beaverbuild 21,447 blocks (40.2%). Hence, Beaverbuild submits proportionally fewer blocks to Agnostic Relay than to other relays. Our primary dataset is, therefore, an unbalanced subsample of all blocks submitted across all relays. We acknowledge this as a limitation of our data.
To explore why builders submit fewer blocks to Agnostic Relay than to other relays, we examine the timestamps associated with the blocks that were shared between Agnostic and at least one other relay. We find that only 5% were received by Agnostic before any other relay. More specifically, among the blocks received by both Agnostic and Ultrasound, just 1% of blocks were received by Agnostic first; between blocks received by both Agnostic and BloXroute only 15% were received by Agnostic first. Agnostic, therefore, has a latency disadvantage relative to the other major relays, which may explain why it receives fewer blocks from builders.
2.2 Transactions
Just as searchers and users pay builders to include their transactions, builders pay proposers by appending a fee reception transaction at the end of each block. In our primary dataset of 15,097 unique blocks, 13,848 (91.7%) contain a fee reception transaction, totaling 12,205 unique fee reception transactions (some blocks differ in content but include the same fee reception transaction). The remaining blocks — those without a fee reception transaction — are low-value blocks in which the builder sets the fee recipient variable directly to the proposer’s address. In these cases, the bid is the sum of payments sent to the fee recipient address. In what follows, we analyze builder bidding behavior separately from users’ and searchers’ activity. To do so, we label “fee reception transactions” as “bids” so that when we refer to “transactions,” we mean transactions that are not fee reception transactions. This leaves us with 10,793 unique transactions excluding bids. Of these, only 5,576 were eventually included in a winning block—either in the cycle in which the transaction first appeared or in a later one.
A notable feature of our primary dataset is that 2,981 transactions — 35% of all unique transactions — originate from just two addresses. The first address,1111110x68d3A973E7272EB388022a5C6518d9b2a2e66fBf. is responsible for 1,873 transactions, of which only 31 were included in a winning block. Notably, only 6 (out of 1,873) transactions from this address appear in blocks not built by Titan, and all 6 are simple token transfers to Binance. The second address,1212120x51C72848c68a965f66FA7a88855F9f7784502a7F submitted 1,108 transactions, with just 97 included in a winning block. Only 249 (out of 1,108) transactions appear in blocks not built by Rsync, and only 4 appear in blocks not built by either Rsync or Titan. These 4 transactions are also simple token transfers, including two to addresses associated with Wintermute, an algorithmic trading fund that operates Rsync Builder. Based on these patterns, we interpret these addresses as searchers integrated with the Titan and Rsync builders, respectively. We refer to them as “Titan-bot” and “Rsync-bot”. [^13] Using a different methodology, Heimbach et al. (2024) also identified Rsync-bot as a non-atomic arbitrage bot, while Titan-bot did not yet exist during the period covered by their study.
To further support our identification of “Rsync-bot” and “Titan-bot,” we analyze whether these searchers submit different transactions to different builders. Using transaction hashes, we find no case in which Rsync-bot sent the same transaction to both Rsync and Titan. However, when comparing transaction logs — i.e., the sequence of operations executed by the transaction and their outputs — we find 203 such cases. The discrepancy between transaction hashes and logs arises because Rsync-bot modifies the fee offered depending on the builder: on average, it pays 18% more to Rsync than to Titan for executing the same transaction. Moreover, we find that transactions sent exclusively to Rsync often include a provision that renders them non-executable if the block’s fee recipient is not Rsync. This behavior further supports our interpretation that Rsync-bot is integrated with Rsync and that the 18% fee differential represents the bot’s targeted profit margin when interacting with non-integrated builders. To perform the same analysis for Titan-bot, we consider winning blocks for our study period that are not in our primary dataset (because they were not submitted via Agnostic relay). Doing so, we find 20 transactions originating from Titan-bot included in blocks won by Beaverbuild. In 3 cases, the same transactions also appear in our dataset in blocks built by Titan. When comparing transaction logs, we find another two transactions in blocks built by Beaverbuild and Titan. In both cases, Titan-bot paid more for inclusion (by 43% and 23%), when the transaction was sent to Titan than Beaverbuild, again, supporting our interpretation that Titan-bot is integrated with Titan.
Almost all transactions from these bots involve token swaps on decentralized exchanges (DEXs): 1,864 for Titan-bot and 1,095 for Rsync-bot. For Titan-bot, the most frequently traded asset pairs are USDC/WETH (334 swaps), USDT/WETH (284 swaps), LINK/WETH (219 swaps), DAI/WETH (202 swaps), and MOGCOIN/WETH (200 swaps). The most-frequently traded asset pairs for Rsync-bot are instead USDT/WETH (315 swaps), USDC/WETH (208 swaps), WBTC/cbBTC (150 swaps), WBTC/WETH (81 swaps), and WBTC/USDT (55 swaps). We later discuss how these bots often trade on the same DEX pools during the same auction cycle. Based on this activity, we infer that both bots likely engage in non-atomic arbitrage — that is, arbitrage between on-chain and off-chain financial markets — and are often in competition.
Swaps on decentralized exchanges (DEXes) are also the most common type of transaction across all unique transactions in our primary dataset. We identify 5,291 swap transactions, of which 3,919 occur on Uniswap V3, 1,228 on Uniswap V2, 142 on Sushiswap, and 101 on Pancakeswap. [^14] We also observe 99 transactions that interact with multiple DEXes, which is why the total number of swaps on individual DEXes exceeds the number of unique swap transactions. Despite their prevalence, only 814 swap transactions — approximately 15% of all swap transactions — were ultimately included on-chain. A small number of addresses, all belonging to searchers, account for the majority of swaps: the top four addresses are responsible for 3,653 swap transactions (about 70%), with Rsync-bot and Titan-bot alone responsible for 2,959 (56%). Swaps are also concentrated among a few pools: the top 11 DEX pools account for 2,872 transactions (54.4%). The most frequently traded token pairs are USDC/WETH, USDT/WETH, MOGCOIN/WETH, and LINK/WETH.
We also identify 807 swap transactions that interact with 42 distinct DEX routers, services that optimize swap execution by splitting trades across multiple liquidity pools and/or using intermediate assets. The most frequently used routers are the Uniswap Universal Router (190 transactions), Uniswap V3 Router (172), 1inch V5 Router (52), and Uniswap V2 Router (42). Of these, 594 transactions were included in a winning block, implying that the majority of swaps included on-chain were routed through a DEX router. The remaining 213 router-based transactions were not included, and the vast majority of these (153) originate from a single user address, 0xf5213a6a2f0890321712520b8048d9886c1a9900. This user submitted 153 transactions calling the Uniswap V3 Router, of which only two were ultimately included in a winning block.
2.3 Transaction delivery
A notable feature of our dataset is that it allows us to distinguish between private transactions shared among multiple builders and exclusive transactions — that is, transactions available to only a single builder. As already discussed, this distinction is important in light of the current concentration in the builders’ market, which is often attributed to builders’ exclusive transactions.
We use data from MempoolGuru to label transactions as either public (i.e., they appeared in the public mempool) or non-public (i.e., they appeared in a block without first appearing in the public mempool). We also incorporate data from MEV Blocker and Flashbots Protect, two private mempools that share transactions with multiple builders (it is important to note that these public sources only classify transactions that were ultimately included in winning blocks).
We find that, of the 5,576 transactions in our primary dataset that were included in a winning block, MempoolGuru classifies 4,124 as public and 1,452 as non-public. Among the non-public transactions, we find that 266 were submitted through MEV Blocker, 40 through Flashbots Protect, and 14 through both — implying that these transactions are private but not exclusive. Using our dataset, we further investigate the remaining 1,132 non-public transactions included on-chain. We find that 642 of them — 45% of all non-public transactions in winning blocks — appear only in blocks proposed by a single builder, indicating that they are exclusive. Of these, 533 appear only in blocks by Titan Builder, 107 only in blocks by Rsync Builder, and 2 only in blocks by builder0x69. In terms of value, among the winning blocks, the sum of all fees paid to block builders for inclusion is 14.67 ETH, of which 12.45 ETH (85%) by exclusive transactions, 1.47 ETH (10%) by private transactions, and 0.75 ETH (5%) by public transactions. Exclusive transactions, therefore, are builders’ dominant source of revenues.
With respect to the 5,217 transactions never included in a winning block, about half (2,853 or 54.7%) are from Titan-bot and Rsync-bot, which, as we already discussed, are exclusive transactions (although, at times, Rsync-bot also sent transactions to Titan). Of the remaining 2,002 transactions, 1,554 (77.6%) transactions are present only in blocks by a single builder and are, therefore, exclusive. Hence, approximately 85% of the non-public transactions that were never included in a winning block are exclusive, roughly double the proportion of exclusive transactions among non-public transactions included in winning blocks.
As already discussed, one limitation of our analysis is that we observe only a small subset of blocks produced by Beaverbuild. As a result, some transactions we classify as “exclusive” may be present in some Beaverbuild blocks that are not included in our dataset. However, in Section 4 we analyze bilateral sharing of transactions between builders and find that while some sharing does occur, it involves relatively few transactions. If we extrapolate this result to transactions shared with Beaverbuild, our estimates of exclusivity should remain approximately accurate. A more important concern is that transactions exclusive to Beaverbuild do not appear in our primary dataset if Beaverbuild did not submit the corresponding block to Agnostic, regardless of whether the transaction was ultimately included in a winning block. This suggests that the share of exclusive transactions among all transactions submitted for inclusion during the study period may be larger than the share of exclusive transactions in our primary dataset.
3 Block 21322649
We now turn to a detailed empirical analysis of a specific auction cycle—namely, the one that resulted in the inclusion of block 21,322,649 on the Ethereum blockchain. This cycle is noteworthy for two reasons. First, bid data across all relays reveals intense competition between Titan Builder and Rsync Builder, with Rsync ultimately winning the auction (see Figures 2(a) and 2(b)). Second, Rsync-bot and Titan-bot actively compete in the same DEX pools, and account for 94%–97% of the total revenue generated by blocks in which they are included. We can therefore closely examine the dynamics of the bidding cycle by analyzing the two bots’ behavior and their respective builders’ bidding strategies.
(a) Builders’ bids across all relays during the auction cycle for Block 21322649
(b) Number of Unique transactions, blocks and swaps vs ETH price change on Binance, during the auction cycle for Block 21322649
Figure 2:
3.1 Description of the data for the auction cycle of block 21,322,649
For this auction cycle, our primary dataset contains 669 blocks and 504 unique transactions, of which 300 (59%) are swaps. Only 22 swaps were included in winning block 21,322,649, 8 from users and 14 from searchers. An additional 8 swaps were included in winning block 21,322,650, 3 more in winning block 21,322,651 and 1 more in winning block 21,322,652. Notably, all the swaps included in these later blocks are from users. A total of 181 transactions are attributed to Titan-bot and 42 to Rsync-bot, all of which are DEX swaps. None of the Titan-bot transactions were included in a winning block. In contrast, 5 transactions from Rsync-bot were included, all in winning block 21,322,649.
(a) Swaps traded volume per block
(b) Percentage of the total DEX volume attributable to Rsync-bot and Titan-bot
(c) Number of Unique Public, Private, and Exclusive transactions
(d) Fees earned from different types of transactions (in ETH)
Figure 3:
As the bidding cycle progresses, we observe a sharp increase in the number of swaps (see Figure 2(b)). This is accompanied by a rise in volume swapped in each block (see Figure 3(a)), as well as in the share of swap volume attributable to the two main searchers, Titan-bot and Rsync-bot (see Figure 3(b)). Together, these two bots account for 94%–97% of the total revenues of blocks in which they are present, indicating that their activity is the primary driver of auction dynamics during this cycle. In contrast to public transactions, which arrive at a relatively uniform rate throughout the bidding cycle, private and, especially, exclusive transactions tend to arrive toward the end of the cycle (see Figure 3(c)). This pattern is also reflected in the distribution of fees paid by different transaction types (see Figure 3(d)).
The fact that high-value transactions tend to arrive later in the bidding cycle should have implications for how builders prioritize them within blocks. For Titan, we identify 248 new transactions that arrived after the submission of their first block in the auction cycle, 219 of which were DEX swaps. Titan included the majority of these late-arriving transactions (over 50%) in the top 1.5% percentile positions within the block, confirming the expectation that they were treated as high-priority or high-value. For Rsync, we identify 59 new transactions, 34 of which were DEX swaps. Surprisingly, the median inclusion position for Rsync’s late-arriving transactions is the 68th percentile, that is, the middle of the block. This pattern holds even for Rsync-bot’s unique transactions, which have a median inclusion percentile of 25%.
(a) Bids by builders
(b) Block revenues by builders
(c) Perc. of block value kept by Titan
(d) Perc. of block value kept by Rsync
Figure 4: Bids and value of blocks for different builders (note the changes in the time scale)
Of the 669 proposed blocks in our primary dataset, 413 were submitted by Titan Builder and 123 by Rsync Builder. These two builders also account for all of the “high-paying” blocks, as shown in Panel (a) of Figure 4. Panel (b) shows that the revenues from the submitted blocks — measured as the sum of payments to the builder — closely track the builders’ bids over time. The remaining panels of Figure 4 examine the portion of block revenues that Titan and Rsync retain versus those paid to the proposer as bids. These panels show that, as the bidding cycle progresses, both builders gradually reduce the share of revenue they retain, eventually approaching zero. The convergence to zero is faster and more pronounced for Rsync (Panel d) than for Titan (see Panel c).
3.2 Similarity of blocks and transaction sharing
(a) 11 seconds
(b) 12 seconds
(c) 13 seconds
(d) 14 seconds
Figure 5: Proportion of transactions shared between blocks by different builders, at different moments in the bidding cycle.
We begin by analyzing whether blocks submitted by different builders differ in their content. Figure 5 shows the fraction of transactions in each builder’s block that also appear in blocks submitted by other builders. These comparisons are made at various points in the bidding cycle. In this analysis, transactions are identified by their hashes.
We observe that, as the bidding cycle progresses, blocks built by different builders become increasingly dissimilar. This divergence is asymmetric, particularly in the case of Titan. Over time, Titan’s blocks contain fewer transactions that are also present in blocks from other builders, while the reverse is not true: an increasing share of transactions in blocks from other builders are also found in Titan’s blocks. A similar, though weaker, pattern is observed for Rsync. One possible explanation is that losing builders share transactions with builders they expect to win the auction. Equivalently, searchers who initially submit to a builder that appears unlikely to win may later resubmit the same transaction to a builder perceived as more competitive. But it is also possible that some builders, particularly those with integrated searchers, receive more exclusive transactions, which tend to arrive later in the bidding cycle.
To distinguish between these two explanations, we examine whether new transactions received by Rsync or Titan previously appeared in blocks by other builders, where a “new transaction” is a transaction included in a block submitted at least one second after a builder’s initial submission, and not present in that first block. We identify three transactions that appeared only in blocks from Flashbots Builder and Buildernet (a Flashbots-affiliated builder) around 10 seconds into the auction cycle. Then, between 12 and 13 seconds in the bidding cycle, these transactions also appeared in blocks built by Titan. None of them were included on-chain. It is plausible that these transactions were submitted via Flashbots Protect, which only logs transactions that are eventually included, and were later shared with other builders once it became clear that Titan and Rsync were prepared to outbid all non-integrated competitors (we provide further evidence of this behavior in the next section). Overall, while transaction sharing does occur, it is limited in scale and does not fully account for the asymmetries observed in Figure 5.
3.3 Disparity in Transaction Execution
Having established that blocks by different builders are different, we now turn to study whether the execution of a transaction changes depending on which builder includes it in a block. For instance, a token swap on a DEX may yield different outcomes depending on the transactions that precede it. We focus on swaps on major DEX protocols — Uniswap V2, Uniswap V3, Pancakeswap, and Sushiswap — that are included in multiple blocks. We compare their execution across blocks using transaction logs, which capture the sequence of contract calls and resulting outputs. This enables us to detect differences in how the same transaction is executed depending on block context.
We begin by examining swap transactions with identical transaction hashes that appear in multiple blocks. As previously discussed, searchers typically modify either the fee or the swap amount during the bidding cycle, resulting in each of their transactions having a unique hash. Consequently, transactions with the same hash across different blocks are likely to originate from users rather than searchers.
The first dimension along which a transaction’s execution may vary depending on the block in which it is included is its execution speed. We identify 12 user swaps that first appeared during the bidding cycle for block 21,322,649 and were still present in the subsequent cycle. Of these 12 delayed transactions, 5 were exclusive to Titan during the 21,322,649 cycle. The remaining 7 were present in blocks by Flashbots, Buildernet, and Titan. [^15] These 7 transactions are different from the 3 transactions identified in the previous subsection because these 3 transactions were present only during the bidding cycle for block 21,322,649. In the following cycle, 3 of the 5 transactions initially exclusive to Titan remain exclusive, while the other 2 change status to “private but not exclusive,” having been submitted via both MEV Blocker and Flashbots Protect. All 9 transactions that were not exclusive to Titan in the second cycle are included in the winning block for that cycle (21,322,650), which was built by Beaverbuild. 2 of the remaining 3 transactions are included in block 21,322,651 (built by Titan), for which we do not have complete bidding data. [^16] The hashes of those transactions are
0x758f12756ba2a91d417a940e311f4500c449e464e0233a4b896203228302af22
0x100825b825744accd39b532d2f7c98c5d8472154b21092294e83f227e9d6d2ff
0x74b6b90aa9ec32adc9e35c15fbd1cf5af37a8ee4d43141785ed21f3bef3f6d8b
The last transaction is a user’s swap that remains exclusive to Titan and included in winning block 21,322,652 (also built by Titan), but fails. [^17] Its hash is 0x74b6b90aa9ec32adc9e35c15fbd1cf5af37a8ee4d43141785ed21f3bef3f6d8b These findings suggest that which builder has access to which transactions affects not only whether transactions are included on-chain, but also when they are included.
The second dimension along which a transaction’s execution may vary depending on the block in which it is included is whether the transaction succeeds or fails. For the auction cycle under consideration, we identify seven transactions that fail in at least one block: six of these fail in all blocks in which they appear, and five are included in the winning block. Among the six transactions that always fail, four also fail when we simulate them at the top of the block in which they were included, suggesting that they were erroneously constructed or misconfigured. At least for this bidding cycle, whether a swap fails does not seem to depend on the block in which it was included.
Finally, the same swap transaction may be executed at different prices depending on the block in which it is included. Among the 12 users’ swaps observed across multiple bidding cycles, only 3 exhibit variation in execution price. Surprisingly, for those 3 swaps, execution quality is better in the later cycles. We also identify 7 users’ transactions with different execution prices across blocks within the bidding cycle for block 21,322,649. Two swap ETH for MogCoin and receive the best execution in a block built by Rsync, where they are preceded by a swap by Rsync-bot selling MogCoin. However, we also find 3 transactions selling MogCoin that would have achieved better execution in blocks built by non-searcher-integrated builders (i.e., neither Rsync nor Titan). In the remaining 2 cases, better execution is also offered by non-integrated builders. Overall, transactions that move in the opposite direction of searchers’ tend to receive better execution in blocks built by searcher-integrated builders, while those swapping in the same direction of searchers’ perform better in blocks built by non-integrated builders.
3.4 Competition between Rsync-bot and Titan-bot.
Having established that the identity of the winning builder affects the quality of users’ transaction execution, we now turn to the behavior of Rsync-bot and Titan-bot. As noted earlier, during this auction cycle, these two bots account for between 94% and 97% of the total value of submitted blocks, making them one of the main drivers of the auction’s outcome.
We identify 42 unique transactions from Rsync-bot, interacting with five DEX pools: USDC/WETH, USDT/WETH, MOGCOIN/WETH (2 pools), and WBTC/cbBTC. For Titan-bot, we observe 181 unique transactions across five DEX pools: USDC/WETH, USDT/WETH, MOGCOIN/WETH (2 pools), and MATIC/WETH. Notably, both Rsync-bot and Titan-bot execute swaps that buy WETH for MOGCoin and buy stablecoins (USDC and USDT) for WETH. This suggests that the two bots are competing for the same arbitrage opportunities between decentralized exchanges (DEXes) and off-chain centralized exchanges.
Our data allow us to study the competition between Rsync-bot and Titan-bot by comparing:
  1. 1.
    the trading volume of each bot and how it evolves over the course of the auction cycle;
  2. 2.
    the fees each bot pays for block inclusion and how those fees change during the auction cycle;
  3. 3.
    the price paid on the DEX, both in terms of the raw execution price (based on token in/out amounts) and the effective price net of the inclusion fee.
Furthermore, because both searchers are integrated with their respective builders, we can interpret their willingness to pay for block inclusion as the expected profits from the arbitrage trade. This, in turn, allows us to infer a “risk-adjusted implied price” on the centralized exchange (CEX) side of the arbitrage path—a component that is typically unobservable. This method provides new insights into the off-chain leg of arbitrage activity.
More precisely, suppose a bot buys a given token for ETH on a DEX at price $p_{\text{DEX}}$ (quoted in ETH) and sells the same token for ETH on a centralized exchange (CEX) at price $p_{\text{CEX}}$. The bot’s profit from the arbitrage trade, in ETH, is:
$\pi=v\cdot(p_{\text{CEX}}-p_{\text{DEX}}),$
where $v$ is the volume of the trade, expressed in units of the non-ETH token. While we do not observe $p_{\text{CEX}}$ directly, we do observe the fee paid for block inclusion, $f$, which we interpret as the bot’s risk-adjusted expected profit. Substituting this into the equation above allows us to compute the implied risk-adjusted CEX price:
$p_{\text{implied}}=p_{\text{DEX}}+\frac{f}{v}.$
If, instead, the trade is the opposite direction—selling a token for ETH on the DEX and buying it on the CEX—the formula becomes:
$p_{\text{implied}}=p_{\text{DEX}}-\frac{f}{v}.$
Note that all prices here are quoted in ETH. In some cases, such as trades involving stablecoins (e.g., USDC or USDT), we may prefer to express prices in terms of the other token. This simply requires inverting the expressions above.
(a) WETH bought on the DEX
(b) Fees paid for inclusion (in ETH)
(c) MogCoin / ETH implied DEX price
Figure 6: Competition between Rsync-bot and Titan-bot on the WETH/MOG Uniswap v3 pool.
(a) Amount of ETH sold on the DEX
(b) Amount of ETH sold on the DEX
(c) Fees paid for inclusion (in ETH)
(d) Fees paid for inclusion (in ETH)
(e) USDT / ETH implied DEX price with Binance price
(f) USDC / ETH implied DEX price with Binance price
Figure 7: Competition between Rsync-bot and Titan-bot on the WETH/USDT and WETH/USDC Uniswap v3 pools.
Figure 6 presents the results of our analysis for MOG/WETH swaps on the Uniswap V3 pool, and Figure 7 replicates the analysis for swaps on the USDT/WETH and USDC/WETH Uniswap V3 pools. The first observation is that, despite the fact that the amount swapped is similar on the three markets, the fee paid by the bots for swapping on the MOG/WETH pool is orders of magnitude larger than that paid for swapping in the other two pools. The MOG/WETH arbitrage opportunity is therefore much more valuable than the other two and drives the value of winning this specific auction cycle. Focusing on the MOG/WETH swaps, we note that Titan-bot maintains a constant trade volume throughout the auction cycle, but its fees increase. We interpret this as evidence that the risk associated with the arbitrage opportunity decreases as the auction progresses, thereby increasing its expected profit. Rsync-bot instead increases both the volume traded on the DEX and its fees as the auction cycle progresses.
Finally, we adjust the base currency in the graphs for the implied centralized exchange (CEX) price so that higher prices are always better for the bots. The main result here is that, for the MOG/WETH arbitrage, Titan-bot’s implied CEX price is 1.9% better than that of Rsync-bot (Note that MOG Coin was not listed on Binance nor Coibase during our study period). For USDT/WETH and USDC/WETH, instead, Rsync-bot seems to have an advantage of approximately 0.17% and 0.02%, respectively. Note also that the implied CEX prices are sometimes above or below Binance’s. Remember that the implied CEX prices are risk-adjusted, that is, the arbitrage risk is reflected in a less favorable implied CEX price. Also, the reported Binance price does not account for trading fees or price impact. These reasons may explain why the bots’ implied CEX prices may be below those on Binance. At the same time, the bots may have sources of liquidity other than Binance or some inventory, which they may use to achieve an even better execution price than Binance’s.
To summarize, we showed that the MOG/WETH arbitrage is significantly more valuable than the other two, and hence, Titan-bot can generate considerably more profits than its competitor. Despite this, Rsync Builder wins the block because, as we already discussed, it increases its bid more aggressively than Titan.
4 Entire Dataset
(a) 11 seconds
(b) 12 seconds
(c) 13 seconds
(d) 14 seconds
Figure 8: Proportion of transactions shared between blocks by different builders, at different moments in the bidding cycle.
We now extend our analysis to all bidding cycles in the dataset. Figure 8 confirms the earlier pattern: blocks proposed by different builders differ in their transaction content, and this difference becomes more pronounced as the bidding cycle progresses.
We also investigate whether transactions are shared among builders. Again, we limit our analysis to transactions shared with Rsync or Titan. For each new transaction added by Rsync or Titan, we check whether it appeared in a block submitted by another builder at least one second earlier. [^18] Again, a builder’s new transactions are those included in a block submitted at least one second after a builder’s initial submission, and not present in that first block. We identify 20 such cases: 3 involving Rsync and 17 involving Titan. Of the 3 Rsync cases, two transactions were initially exclusive to Beaverbuild and were later included in a block by Rsync, which ultimately won. The third, a searcher’s swap, appeared in blocks from five different builders, but only entered a Rsync block four seconds after its first appearance (again, Rsync won the block). Of the 17 transactions later included by Titan, 8 were submitted via the Flashbots private mempool. Five originated from a single searcher address, 0x89a99a0a17d37419f99cf8dc5ffa578f3cdb58b5, and appear in Flashbots Protect’s dataset as shared with more than 15 builders. These transactions remained exclusive to Flashbots Builder for approximately 2.6 seconds before being included in Titan’s blocks. Another five transactions from the same address follow the same pattern (initial exclusivity to Flashbots followed by appearance in Titan’s blocks) but were not included on-chain. It is plausible that they were also shared via Flashbots Protect but were excluded from the dataset, which only logs transactions that are eventually included.
In summary, transaction sharing between builders during the bidding cycle does occur, but is relatively rare. In most cases, it involves transactions shared via Flashbots Protect and exclusive to Flashbots builder during the initial part of the bidding cycle.
4.1 Time of inclusion in a winning block
We identified 1,288 transactions that appeared in multiple bidding cycles. These transactions appeared in a bidding cycle before the winning block was selected, were not included in the winning block, and appeared again in the subsequent bidding cycle. Among these, 871 were present for two cycles, 198 for three cycles, and 108 for four cycles before being included in a winning block. Remarkably, one transaction remained present for 2,437 bidding cycles before inclusion; it was exclusive to the builder payload.de and was ultimately included only when payload.de won. In addition, 28 transactions appeared in multiple cycles but were never included in a winning block.
Among the transactions present in multiple bidding cycles, only 17 are searchers’ transactions. Of the remaining transactions, 701 first appeared in our dataset as public, 189 as private, and 381 as exclusive (with 374 exclusive to Titan Builder). inlineinlinetodo: inlineMaybe for another paper, but why are so many public transactions not picked up? Is the public mempool being ignored? Have private mempools crowded out public mempools? Has EIP -1559 failed, meaning: it is not enough to pay base fee + something to be included, but, again, users need to strategize on how much to pay? Transactions that are initially public remain public throughout all subsequent bidding cycles, and very few (13) transactions that begin as private later appear in the public mempool. However, the majority of exclusive transactions present over multiple bidding cycles change status over time. Specifically, 162 exclusive transactions remain exclusive across all cycles in which they appear, while 174 exclusive transactions transition to private.  [^19] We find a further 33 transactions that were initially exclusive to a builder and were included in a later auction cycle by that same builder. However, the auction cycle in which these transactions were included is not in our dataset. Hence, we cannot determine whether these transactions remained exclusive to the builder or transitioned to private. We also identify 12 transactions that were initially exclusive but later appeared in the public mempool. Of these, 9 were present during two bidding cycles, and 3 were present during three cycles. All are user transactions: three interacted with a DEX via the 0x router, and four via the 1inch router. [^20] These transactions are: 0x1daea1584d684385fb25209bad3a49c54a1e2c500e27543334c7b3e695a12ffb, 0x6605c1c36fcd3d696844519d6ff7199fe4ddd34fb4f3e0ed4da71aedfcf1c633, 0x3dadbab8e85b6e48a2c33f9029ea4ebebdd3def8b9d907faff15ef81527de968, 0x9c7364eddb94fe659bc8104585ff86a8a234480fe00a9f9aa3eb410e572ebdf2, 0x9d476dbc524651cbe26bcafab99f6fe40a823a5588fa5cd5002e5b414a8a2f8c, 0x417440b083bdcb529663c6615c32d89a92beb7863f1a5d6b12bca5b94f4a0f87, 0xe56caee27e79dee8a4c7ed6f945ccbc2cc733d761e19bd68329dc364c61e73a7, 0xae5b88bc0bffa04811eda0dfdcd7224599e2757339e18b2d57f4f40d61902a4c, 0xd41139b67a193045b4d60266c084f6dadd0cb95dccd7b2a06ea68653d6a313ec, 0x597fc0cb0610d1e3ea2baf0f6f90d0d741a6429bfdd9b375bfd2493b70d41dcd, 0x44e8105d5c672195877dc1ed970511e0f0a15fc25cbcafe71d51d428d54c215d, 0xd74718e6d8d58bc33334c0c2f6642d9dec49496329329d8fdcb1e698e81abc94.
To summarize: the vast majority of searchers’ transactions appear in only a single bidding cycle and are either included in the winning block or discarded. Hence, transactions present across multiple bidding cycles are predominantly users’ transactions. Given that our dataset includes 5,947 user transactions, we estimate that approximately 21% users’ transactions are delayed. This delay is partly explained by the fact that 30% of these transactions are, at least initially, exclusive to a single builder. This is surprising, as users should submit transactions privately to multiple builders or even broadcast them publicly (in case of simple transfers). Even more startling is that some initially exclusive transactions appear in the public mempool after one or two bidding cycles. These transactions are first delayed and then exposed to various attacks, which is clearly not in the users’ best interest.
4.2 Successful execution
A transaction may succeed or fail depending on the block in which it is included. To explore this possibility, we consider each transaction hash/block combination and classify it as a fail if its simulation outcome is status = 0, or if it produces no transaction log despite the same transaction generating one in another block. A success is defined as any transaction/block pair not classified as a failure. Under this definition, “failures” include both outright failures and reverts, such as a swap reverting because its execution price would be below the user’s specified worst acceptable price. In contrast, a “success” indicates that the transaction would have executed as intended, had the corresponding block been added to the chain. Among the 49 transactions that sometimes fail and sometimes succeed, 44 are swaps. These account for 318,742 transaction-block observations. We therefore focus our analysis exclusively on swaps.
We create a binary (dummy) variable equal to 1 if a transaction executes successfully, and 0 otherwise. We then run a series of regressions using the following explanatory variables:
  • Time since Last Block: the time elapsed since the previous block, measured at the moment the block was submitted for inclusion.
  • Tx index: the position of the transaction within the block, where lower values indicate earlier placement (and thus earlier execution). For example, Tx index = 1 indicates the first transaction in the block, Tx index = 2 the second, and so on. [^21] We also tested a normalized specification in which transaction position is scaled from 1 to 100, with 1 indicating the first and 100 the last transaction in the block. All results are robust to this alternative.
  • A dummy variable indicating whether the block was built by Titan.
  • A dummy variable indicating whether the block was built by Rsync.
  • A dummy variable indicating whether the block contains a transaction from Titan-bot.
  • A dummy variable indicating whether the block contains a transaction from Rsync-bot.
We also include transaction fixed effects: that is, a dummy variable for each unique transaction hash. These control for transaction-specific characteristics and capture the average likelihood that a transaction executes successfully across all blocks in which it appears. The coefficients on the other variables then measure how execution probability varies depending on block timing, builder identity, transaction ordering, and the presence of searcher bots.
Finally, we construct two subsamples: swaps in the same direction as those by Titan-bot and Rsync-bot, and swaps in the opposite direction. To do so, for each swap by either Titan-bot or Rsync-bot, we record the target pool, its direction, and the auction cycle (noting that, in our data, all bots’ swaps on a given pool and during the same cycle are in the same direction). We then identify all other swaps occurring in those same pools and cycles, and divide them between those “in the same direction” as the bots and those “in the opposite direction” as the bots.
Table 1 reports three regressions, each on a different samples. The first regression yields mixed results: being included in a block built by Titan or Rsync reduces the probability of successful execution, but the presence of transactions from the corresponding bot increases it. In contrast, the second and third regressions reveal a clearer pattern. Swaps in the same direction as the bots are significantly more likely to fail when they are included in blocks built by Rsync or Titan, especially if those blocks also contain transactions from the bots themselves. The opposite holds for swaps in the opposite direction: their likelihood of success increases under the same conditions, although in this case the main effect is due to the presence of the bots. We also observe that swaps included later in a block (i.e., with a higher transaction index) are slightly less likely to fail. This result is, however, difficult to interpret, as transaction ordering within a block is determined by the builder. Finally, transactions are more likely to fail when included in blocks submitted later in the bidding cycle.
Using the estimated coefficients, we perform back-of-the-envelope calculations to quantify the effects. Compared to inclusion in a block built by a non-searcher-integrated builder, a swap in the same direction as Titan-bot or Rsync-bot is approximately 18% less likely to execute successfully when included in a block built by Titan or Rsync that also contains a transaction from the corresponding bot. By contrast, a swap in the opposite direction is approximately 1% more likely to execute successfully under the same conditions.
Table 1: Probability of Success
All Swaps Same Dir. as Bots Opp. Dir. as Bots
Time Since Last Block -0.0038*** -0.0217*** -0.0012***
(0.0001) (0.001) (0.000)
Tx Index 0.0001*** 0.0006*** 5.74e-06
(0.0000) (3.18e-05) (9.23e-06)
Is Titan Builder -0.0078*** -0.1341*** -0.0144***
(0.0008) (0.005) (0.001)
Is Rsync Builder -0.0148*** -0.1592*** -0.0021
(0.0006) (0.006) (0.002)
Has Titan-bot tx 0.0024*** -0.0423*** 0.0286***
(0.0005) (0.004) (0.001)
Has Rsync-bot tx 0.0075*** -0.0301*** 0.0095***
(0.0005) (0.004) (0.001)
Observations 318742 17224 26705
R-squared 0.744 0.556 0.493
Tx fixed effect yes yes yes
    $p<0.1$,
    $p<0.05$,
    $p<0.01$
Note: Standard errors in parentheses. Column titles refer to subsample type.
4.3 Execution price
The way a swap is included in a block may also influence its execution price. To investigate this possibility, we examine all swap in blocks where they executed successfully and the execution price can be computed. This yields 272 unique transaction hashes, 112,747 transaction-block observations, and 342,528 swap-block observations, which will be our unit of analysis (note that many transactions perform multiple swaps). inlineinlinetodo: inlineIt is kind of interesting that the majority of transactions performing swaps perform more than 1 swap
For each swap, we compute its average execution price across all blocks, denoted by $p_{\text{avg}}$. We then calculate, for each transaction-block observation, the percentage deviation from this average:
$p_{\text{norm}}=\frac{p-p_{\text{avg}}}{p_{\text{avg}}}\times 100$
where $p$ is the execution price of the swap in a given block. We adjust the base currency so that higher values of $p_{\text{norm}}$ always reflect better execution from the user’s perspective. The distribution of $p_{\text{norm}}$ has a median of $-1.41\times 10^{-14}$, with a minimum of $-10.53$ and a maximum of $13.47$ basis points.
Table 2 reports the coefficients from three regressions of $p_{\text{norm}}$ on the same set of explanatory variables and subsamples used in the success/failure analysis. As with the previous table, the regression on the full sample yields ambiguous results: being included in a block built by Titan is associated with slightly worse execution prices, while the presence of a transaction from Titan-bot in the same block improves execution. For Rsync, the pattern is reversed. These opposing effects largely offset each other. In contrast, the regressions restricted to swaps in the same direction as either bot show a clearer trend: execution prices are between 9 and 12 basis points worse when such swaps are included in blocks built by Rsync or Titan (the presence of the two bots has a negligible effect). For swaps in the opposite direction from the bots, results align with those from the success/failure regressions: being included in a block built by Rsync or Titan improves the execution price by between 9 and 27 basis points. The effect remains positive (but smaller) if the block also contains a Rsync-bot or Titan-bot transaction. Additionally, being placed later in the block (i.e., with a higher transaction index) has a small but consistently positive effect on price. Finally, while many coefficients are statistically significant, the regressions explain only a modest share of the variation in $p_{\text{norm}}$, as reflected in the low $R^{2}$ values.
Table 2: Normalized price $p_{norm}$
All Swaps Same Dir. as Bots Opp. Dir. as Bots
Time since Last Block -0.0022*** -0.0430*** 0.0163***
(0.0005) (0.002) (0.002)
Tx Index 0.0003*** 0.0001*** 0.0007***
(0.0000) (4.01e-05) (4.28e-05)
Is titan builder -0.0053** -0.1206*** 0.0923***
(0.0024) (0.006) (0.007)
Is rsync builder 0.0198*** -0.0924*** 0.2758***
(0.0031) (0.008) (0.009)
Has titan-bot tx 0.0048** 0.0068* -0.0640***
(0.0019) (0.005) (0.006)
Has rsync-bot tx -0.0235*** -0.0038 -0.0402***
(0.0019) (0.005) (0.006)
Observations 342528 15404 25718
R-squared 0.002 0.142 0.073
Tx fixed effect yes yes yes
    $p<0.1$,
    $p<0.05$,
    $p<0.01$
Note: Standard errors in parentheses. Column headers refer to different subsamples.
inlineinlinetodo: inlineAt some point we should try to increase the R2 of the regression - that is, we should find out what other variable has explanatory power on $p_{norm}$
4.4 Competition between Searchers
In the previous section, we highlighted the importance of the competition between Rsync-bot and Titan-bot. It turns out that these two searchers compete intensely not only during the auction cycle of winning block 21322649, but throughout our entire dataset. inlineinlinetodo: inlineThere is a bit of a gap in our reasoning here: in the previous section, we showed that those two bots are responsible for the vast majority of the value of the blocks, which is why we study them. Can we compute the same for the entire dataset? Are these two bots ”important” also when looking at all bidding cycles?
We identify 57 instances in which Titan-bot and Rsync-bot compete for execution in the same DEX pool during the same auction cycle. In 52 cases, one of the tokens swapped is WETH, allowing us to apply the same methodology introduced in the previous section — namely, deriving an implied risk-adjusted CEX price for each bot. To enable direct comparison with observed market data, we further restrict the sample to tokens traded on Binance. This yields 29 auction cycle/DEX pool combinations for which we can compare the bots’ implied prices to contemporaneous Binance prices. We plot the implied DEX price alongside the Binance price for all 29 cases in Appendix 0.A. Note that Binance reports prices at one-second intervals. As expected, the implied DEX prices generally track the Binance price closely. This correlation is particularly strong in cases involving the USDT/WETH and USDC/WETH pairs, which account for the majority of the sample. A weaker but still visible relationship is observed for the UNI/WETH and AAVE/WETH pairs. In contrast, for the DAI/WETH pair, the implied DEX and Binance prices do not exhibit a consistent relationship, likely because Binance is not the primary centralized exchange for DAI trading. This is confirmed by the fact that the total volume on the DAI/ETH market on Binance during our study period was only 2.5 ETH. [^22] For comparison, total volume on Binance during the study period for the other markets was: 1,487 ETH for the USDC/ETH market, 16,209 ETH for the USDT/ETH market, 1,374 ETH for the WBTC/ETH market, 6.9 ETH for the UNI/ETH market, 36 ETH for the AAVE/ETH Binance market.
We complement the visual analysis with a regression approach. For each second $s$, we compute the average implied CEX price and trade volume for Rsync-bot and Titan-bot by aggregating all trades executed by each bot between seconds $s-1$ and $s$. We denote the average implied CEX price by $p_{\text{implied}}$, and associate each observation with timestamp $s$11todo: 1It may be more accurate to center the timestamp at $s-0.5$ or $s+0.5$; we leave this refinement for future work. For technical reasons, we restrict our analysis to the USDT/WETH and USDC/WETH trading pairs. [^23] As discussed earlier, Binance prices for some token pairs appear unreliable. More importantly, the number of observations, values of $p_{\text{implied}}$, and trade volumes vary considerably across tokens. We have 104 observations for WETH/USDC and WETH/USDT combined, but only 24 for WETH/DAI, 18 for WETH/UNI, and 5 or fewer for WBTC/WETH and AAVE/WETH. Moreover, the average value of $p_{\text{implied}}$ is around 3,500 for stablecoin pairs (USDC, USDT, DAI), but drops to 261 for UNI/WETH, 15 for AAVE/WETH, and 0.037 for WBTC/WETH. Trade volumes also vary: average volume is 320.3 ETH for WETH/WBTC, compared to 43.1 ETH for other assets. As a result, separate regressions on the less common pairs are underpowered, while pooled regressions suffer from multicollinearity due to token-specific heterogeneity. This filtering results in 104 total observations: 53 for Rsync-bot and 51 for Titan-bot. We regress $p_{\text{implied}}$ on the contemporaneous Binance price, as well as its one-second lag and lead. The regression results, shown in Table 3, indicate that only the contemporaneous and one-second lagged Binance prices are significantly correlated with the bots’ implied CEX prices. Among these, the contemporaneous price has the strongest explanatory power. Therefore, the contemporaneous Binance price serves as the most appropriate benchmark for evaluating $p_{\text{implied}}$.
Table 3: Implied Price on Different Binance Prices
Coefficient Standard Error
Binance Price (+0 sec.) $0.6587$*** $0.106$
Binance Price (+1 sec.) $-0.1197$ $0.128$
Binance Price (+2 sec.) $-0.1306$ $0.095$
Binance Price (-1 sec.) $0.4783$*** $0.097$
Binance Price (-2 sec.) $0.1001$ $0.060$
Constant $46.192$* $19.030$
Observations 104
R-squared 0.997
    $p<0.1$,
    $p<0.05$,
    $p<0.01$
We then proceed to compare each observation of $p_{\text{implied}}$ to the corresponding contemporaneous Binance price, $p_{\text{binance}}$. To interpret the execution quality from the bot’s perspective, we distinguish between two cases. If a bot sells ETH on-chain, it must be buying ETH on the CEX; in this case, a lower CEX price is better. We therefore define
$p_{\text{diff}}\equiv p_{\text{binance}}-p_{\text{implied}}.$
Conversely, if the bot buys ETH on-chain we define
$p_{\text{diff}}\equiv p_{\text{implied}}-p_{\text{binance}}.$
In both cases, a higher $p_{\text{diff}}$ indicates better execution than Binance. We then calculate
$p_{\text{Improvement}}\equiv\frac{p_{\text{diff}}}{p_{\text{binance}}}\times 100$
as the percentage price improvement relative to the contemporaneous Binance price. The minimum, maximum, and average values of $p_{\text{Improvement}}$ are -0.017652, 0.064844, and 0.011934, respectively.
Table 4 reports the coefficients of two separate regressions. Both include a constant, a dummy variable equal to 1 if the observation is from Titan bot, and time measured in seconds since the last block. The difference between the two regressions is that the second includes swap volume (in WETH) in piecewise linear form, where each piece corresponds to 50 ETH. The coefficient on the constant represents the average percentage price improvement ($p_{\text{Improvement}}$) for a zero-volume swap by Rsync-bot, assuming the block was built at time zero in the bidding cycle. The coefficient on Is Titan bot captures the difference in execution quality between Titan-bot and Rsync-bot. The results indicate that the bots achieve better execution than the Binance price, and that Rsync-bot outperforms Titan-bot. Execution quality seems to decline with trade size, although this result needs to be taken with a grain of salt because it is driven by the only 2 trades of volume larger than 150. Using the estimated coefficients, we perform back-of-the-envelope calculations. For a swap of volume lower than 150 ETH the implied CEX price for ETH is approximately 4.2 basis points better than the contemporaneous Binance price for Rsync-bot, and 3.4 basis points better for Titan-bot.
Table 4: Regression of percentage price improvement over Binance ($p_{\text{Improvement}}$) with and without piecewise volume
(1) (2)
Time since Last Block -0.0015 -0.0019
0.001 0.001
Is Titan Bot -0.0085*** -0.0081***
0.003 0.003
Volume X (Volume lower than 50) -0.0015
0.004
Volume X (Volume between 50 and 100) -0.0043
0.004
Volume X (Volume between 100 and 150) -0.0061
0.004
Volume X (Volume greater than 150) -0.0417***
0.009
Constant 0.0368** 0.0420***
0.018 0.013
Observations 104 104
R-squared 0.088 0.272
    $p<0.1$,
    $p<0.05$,
    $p<0.01$
Note: standard errors in parentheses
5 Conclusions
We study the creation of the Ethereum blockchain by analyzing proposed blocks that were ultimately discarded. To the best of our knowledge, this is the first systematic analysis of non-winning blocks. Our findings illuminate key aspects of Ethereum’s block-building process, with implications for ongoing policy discussions and the academic literature.
While informative, our results should be interpreted in light of certain limitations. As noted, the dataset underrepresents some major builders (notably Beaverbuild) and spans only eight minutes. Moreover, our analysis is primarily descriptive and does not identify the economic mechanisms driving the observed patterns. To address these gaps, we hope that relays, builders, and other participants in the transaction supply chain will make more comprehensive data available in the future, enabling a deeper and more rigorous understanding of how Ethereum is made.
Appendix 0.A Competition between Rsync-bot and Titan-bot on the same pool during the same slot
(a) USDC Pool 1
(b) USDC Pool 2
(c) USDT Pool 1
(d) USDT Pool2
Figure 9: Block 21322626
(a) USDC Pool 1
(b) USDT Pool 1
(c) USDT Pool 2
Figure 10: Block 21322630
(a) USDC Pool 1
(b) USDC Pool 2
Figure 11: Block 21322631
(a) USDC Pool 1
(b) USDC Pool 2
(c) USDT Pool 1
Figure 12: Block 21322648
(a) USDC Pool 1
(b) USDT Pool 1
Figure 13: Block 21322649
(a) USDC Pool 1 Block 21322657
(b) USDC Pool 1 Block 21322658
(c) USDT Pool 2 Block 21322659
(d) USDT Pool 3 Block 21322660
Figure 14: Block 21322657 - 21322660
(a) DAI Block 21322635
(b) DAI Block 21322659
(c) DAI Block 21322660
(d) DAI Block 21322660
(e) DAI Block 21322660
Figure 15: DAI all blocks
(a) UNI Block 21322628
(b) UNI Block 21322639
(c) UNI Block 21322643
(d) UNI Block 21322643
(e) UNI Block 21322653
(f) UNI Block 21322653
Figure 16: UNI all blocks. Note that panels (e) and (f) plot the same data but at different time scales.
Figure 17: AAVE Block 21322658
Figure 18: WBTC Block 21322641