### v3

Abstract
Blockchain’s economic value lies in enabling financial and economic transactions that do not require trusted, centralized intermediaries. In practice, however, transactions must pass through several intermediaries before being included on-chain. We study empirically whether this process undermines blockchain’s stated benefits by assembling a novel dataset of 15,097 non-winning Ethereum blocks—blocks proposed by builders but not ultimately selected for inclusion. We show that 21% of user transactions are delayed: although proposed in some candidate blocks, they are not included in the winning block. Approximately 30% of these delayed transactions are exclusive to a single losing builder, indicating that transaction routing materially affect inclusion outcomes. We further document substantial heterogeneity in execution quality: both the probability of successful execution and the execution price of users’ swaps vary across candidate blocks. Finally, we study two arbitrage bots trading between decentralized (DEX) and centralized exchanges (CEX). We document intense competition for the same arbitrage opportunities and estimate that these bots trade USDC/WETH and USDT/WETH on centralized exchanges at prices approximately 2.8 basis points more favorable than contemporaneous Binance prices.
Keywords: Ethereum, Proposer-Builder Separation (PBS), Order Flow.
1 Introduction
Blockchains are often described as “immutable ledgers” because, once a transaction is included in the canonical chain, it triggers a deterministic sequence of irreversible state changes. This immutability is central to blockchain’s economic value, as it enables financial and economic transactions that do not rely on trusted, centralized intermediaries.
This static view, however, obscures the probabilistic and competitive process that determines which transactions become immutable. In Ethereum, transactions are generated either by users or by searchers—automated bots that create transactions in response to, for example, price movements in financial markets. Users and searchers submit transactions to builders through several channels: the public mempool (a publicly observable repository of pending transactions), direct bilateral communication, or intermediaries known as private mempools (or private RPCs). Builders aggregate and order these transactions into candidate blocks and submit them to relays, each of which forwards a single block to the proposer, who ultimately appends the next block to the chain. [^1] See Section 1.1 for a detailed description of this process, commonly referred to as Proposer–Builder Separation (PBS). The overall process resembles an auction aimed at selecting the block containing the most valuable transactions, as determined by their fees. This auction, however, is highly noisy: the value of including a transaction often depends on rapid price movements in financial markets; builders continuously construct and resubmit blocks; and communication latency plays a central role in determining the final outcome. As a result, for each block ultimately added to the canonical chain, up to 5,000 candidate blocks are constructed, propagated, and then discarded together with their transactions.
An important research question is therefore whether the intermediaries that control transaction inclusion undermine blockchain’s promise of disintermediation. Addressing this question empirically requires comparing discarded Ethereum blocks both with one another and with the block that ultimately appears on-chain. If, for example, an intermediary’s behavior causes a user transaction to appear only in a subset of candidate blocks, access to the blockchain may be delayed or denied because the winning block may omit that transaction. More subtly, transaction outcomes—particularly for swaps—are far less deterministic once we recognize that the same transaction may be included differently across candidate blocks, either by occupying different positions or by being bundled with different other transactions. At the same time, it is also possible that the process of block construction and selection has little impact on ordinary users if all proposed blocks are sufficiently similar from their perspective. Despite the importance of this question, empirical analysis has until recently been infeasible because discarded blocks are not publicly disclosed.
In this paper, we overcome this limitation by assembling novel dataset of non-winning blocks submitted between Ethereum blocks 21,322,622 and 21,322,660 (39 winning blocks), roughly from 2:37:35 PM to 2:45:25 PM UTC on December 3, 2024. The dataset is approximately 20 GB in size and contains 15,097 blocks, 10,793 unique transactions (identified by their hashes), and 2,380,014 transaction–block pairs. [^2] The dataset includes all blocks submitted during the study period through a relay called Agnostic. See Section 2 for a detailed description of the data. We simulate each candidate block to recover transaction-level execution outcomes, and we complement these data with several publicly available sources, including second-by-second price data from Binance.
Our main finding is that the behavior of intermediaries controlling transaction inclusion generates widespread delays. Approximately 21% of users’ transactions appear in some proposed blocks but not in the corresponding winning block, postponing on-chain inclusion by one or more blocks. [^3] We distinguish between users and searchers (see Section 2.2 for a description of the heuristic used to make this distinction). Searchers’ transactions are almost never delayed: they are either included immediately or they are discarded. Strikingly, about 30% of these delayed user transactions appear exclusively in blocks submitted by a single losing builder. This pattern is difficult to rationalize, as users do not benefit from sharing transactions exclusively with a single builder rather than privately with multiple builders. We also identify 12 user transactions that remain exclusive to a losing builder for two or three consecutive auctions before later appearing in the public mempool. These transactions are not only delayed but also subsequently exposed to potential attacks; one is in fact subject to a sandwich attack. In four cases, the same builder that initially had exclusive access ultimately includes the transaction on-chain. We speculate that a private mempool operator initially shares transactions exclusively with a single builder and reverts to the public mempool when that builder fails to win successive auctions.
We also find that the process of block construction and selection introduces substantial randomness into execution outcomes, with execution risk concentrated among a small subset of transactions. To establish this result, we consider each user swap in our data and all blocks in which such swap appears. We simulate those blocks to determine the swap’s execution had that block been included on-chain. We then compute the fraction of blocks in which the swap executes successfully, and average and standard deviation of its execution prices across all blocks in which it appears (conditional on successful execution). We find that the vast majority of users’ swaps (96.7%) execute successfully in all blocks in which they appear. However, a small fraction fail very frequently: 1.7% fail in at least half of the blocks in which they appear, and 1.1% fail in more than 80% of such blocks. Execution prices are similarly skewed. For half of the swaps in our sample, execution prices are quite concentrated: at least two-thirds of realizations deviate by less than 4.3 basis points from the swap’s average execution price across blocks. By contrast, for 7% of swaps execution prices are highly dispersed, with at least one-third of realizations deviating by more than 100 basis points from the swap’s average.
Regression results show that a nontrivial share of the variation in both execution success and execution prices can be explained by the presence of other transactions in the same block. Restricting the sample to swaps that sometimes fail and including transaction fixed effects, we find that swaps trading the same asset pair and in the same direction as the two main arbitrage bots in our data (see the next paragraph) are approximately 18% less likely to execute when included in a block that also contains transactions from those bots. By contrast, swaps trading the same asset pair but in the opposite direction are about 1% more likely to execute under the same conditions. A similar, though quantitatively smaller, pattern emerges for prices: swaps in the opposite direction of the bots trade at more favorable prices when included in blocks that also contain bot transactions, whereas swaps in the same direction trade at worse prices.
We then turn to the determinants of which block wins. To do so, we exploit a salient feature of our dataset: roughly one-third of all transactions originate from just two addresses that almost exclusively perform swaps. We identify these addresses as competing arbitrage bots. They frequently trade the same asset pairs in the same direction; they frequently replace their transactions within an auction cycle; and, when they compete for the same arbitrage opportunity, their swaps are mutually exclusive, in the sense that no proposed block contains transactions from both bots. These bots’ transactions are also highly valuable, accounting on average for about 60% of the total transaction fees in the blocks in which they appear. Their competition therefore plays a central role in determining auction outcomes.
We study this competition by simulating each block that contains a swap transaction from either bot and extracting the associated swap volume, on-chain execution price, and transaction fee. If we further assume that each bot’s fee equals its expected profit from inclusion on-chain, we can infer an implied centralized exchange (CEX) price—that is, the price of the off-chain leg of the arbitrage trade—at different points in time, corresponding to the block construction timestamp. [^4] This assumption is justified by the fact that each bot appears to be integrated with or exclusive to a different builder (see Section 2.2). If the assumption is relaxed, the bot’s fee represents only a lower bound on expected profits, implying that the implied CEX prices we recover may be even more favorable to the bots than our estimates suggest. We find that when the bots compete to rebalance either a USDT/WETH or a USDC/WETH pool, their implied CEX prices are nearly identical, indicating intense competition. Moreover, these implied prices are approximately 2.8 basis points better than contemporaneous Binance prices for both bots.
1.1 Background: the journey of an Ethereum transaction.
In Ethereum, the right to add the next block to the blockchain is assigned to a proposer, randomly selected from validators who have staked at least 32 ETH. In the vast majority of cases (over 95%), the proposer delegates block construction to professional entities known as builders. Builders collect transactions from both the public mempool and private channels and use sophisticated algorithms to determine which transactions to include in a block and in what order. This arrangement is known as Proposer–Builder Separation (PBS).
The relation between builders and proposers is facilitated by relays. Their primary role is to prevent proposers from altering the submitted blocks. [^5] Technically, this is achieved by requiring proposers to sign an empty block, which is then completed by the relay. For more details, see https://docs.flashbots.net/flashbots-mev-boost/relay. Once a builder creates a block, it submits it to a relay along with a proposed payment to the proposer. The relay then forwards the highest-paying block and its corresponding payment to the proposer upon request. This process resembles an ascending price auction because the relay continuously broadcasts the value of the highest-paying bid—though this broadcast is observed with a delay due to network latency. An auction cycle lasts approximately 12 seconds, after which the winning block is selected, and a new auction begins. As already mentioned, latency plays a crucial role, as proposers are geographically dispersed, and arbitrage transactions often depend on rapid price movements in traditional “off-chain” financial markets. To account for geographic dispersion, multiple relays operate worldwide, with builders submitting to several at once. Our dataset, for instance, comes from a relay called Agnostic Relay.
Users can submit transactions through several channels. The default option is to broadcast them to the public mempool, a publicly observable repository of pending transactions maintained collectively by Ethereum nodes. Broadcasting transactions to the public mempool, however, exposes users to attacks. For example, when a user submits a swap on an automated market maker (AMM), a malicious searcher may mount a sandwich attack by front-running the swap with the same trade and subsequently back-running it with the opposite trade, thereby “buying low” and “selling high” at the user’s expense. To mitigate this risk, many users rely on private mempools—services that allow transactions to be sent directly to one or more builders without being broadcast to the public mempool. During the study period, five major private mempool operators were active. Four of them—MEV Blocker, Flashbots Protect, Merkel, and Blink—state in their documentation that they privately share transactions with multiple builders (see janicot2025privatemevprotectionrpcs for a detailed comparison). By contrast, the fifth operator, MetaMask Smart Transactions, does not disclose its transaction-sharing policy.
The choice between submitting a transaction via the public mempool or through one of the private mempools is typically determined by the user’s wallet, a software application that automates transaction creation and broadcasting. Wallets generally provide a default option that users can modify—although this is not always the case, especially for mobile wallets. For example, by default, users of the popular MetaMask wallet submit Ethereum transactions via MetaMask Smart Transactions.
1.2 Relevant literature
Ethereum’s process of block creation, submission, and selection has been the subject of considerable debate (for a summary of this debate, with a particular emphasis on the economic aspects, see JOHN2025102718). Our results contribute directly to these discussions.
One of the main criticism of PBS is that it enables searchers to efficiently extract value from users. Quantitatively, the main concern is arbitrageurs profiting from price discrepancies between decentralized exchanges (DEXes) and centralized exchanges (CEXes). The resulting loss to on-chain liquidity providers is referred to as Loss-vs-Rebalancing (LVR), a concept introduced by milionis2022automated. Subsequent studies have attempted to identify arbitrageurs (heimbach2024non, wu2025measuring) and quantify LVR (canidio2023arbitrageurs; fritsch2024measuring). These analyses, however, must rely on strong assumptions because on-chain data alone do not reveals whether a given swap was part of an arbitrage strategy. Furthermore, the prices at which arbitrageurs trade on centralized venues are usually unobservable.
Our dataset overcomes these limitations in two ways. First, we can identify arbitrage bots by observing their competition for the same arbitrage opportunities. In this respect, our approach resembles that of aquilina2022quantifying, who identify high-frequency trading races in traditional financial markets by analyzing all messages received by the London Stock Exchange over a 9 week period. Their rationale is that when multiple arbitrageurs try to exploit the same arbitrage opportunity, only one transaction will succeed, and all others will fail. Similarly, we observe multiple arbitrage bots trying to exploit the same opportunity by submitting transactions, with only one of them eventually successfully included on-chain. Second, by focusing on builder-integrated searchers engaging in CEX-DEX arbitrage, we can infer their expected profits and thus the price at which they trade on centralized exchanges.
An important step in our analysis is establishing that our two main arbitrage bots are integrated with, or exclusive to, a different builder—one with a builder called Titan and the other with a builder called Rsync. Our analysis therefore relates to existing studies of the relationship between searchers and builders, in particular pai2024structural and wu2025measuring. pai2024structural develop a theory of vertical integration between builders and searchers, in which an integrated searcher’s payment to the builder reflects the full value of including its transaction. wu2025measuring use on-chain data to identify searchers engaging in CEX–DEX arbitrage and study their interactions with major builders; they also identify the two bots we analyze as integrated with, or exclusive to, Titan and Rsync, respectively.
capponi2024price also studies the competition between arbitrage bots for the same CEX-DEX arbitrage opportunities. However, they consider an earlier period in which searchers competed by submitting transactions to the public mempool and engaging in priority fee auctions — i.e., because all transactions are included in the block and ordered by priority fee, the winning searcher is the one paying the highest fee. The current market structure and nature of the competition between searchers are very different, with searchers submitting directly to builders, builders selecting which searchers’ transactions to include, and frequent integration between searchers and builders. [^6] See also capponi2024price-private, who develop a theory of searchers’ competition when transactions are sent privately to builders.
Existing work studying the strategic behavior of intermediaries responsible for transaction inclusion has largely focused on theoretical analysis. lehar2023battle and CAPPONI2025104132 study the trade-off between submitting transactions via the public mempool and via private channels, and characterize the resulting equilibrium sorting of transactions across submission routes; CAPPONI2025104132 also provides empirical support for its model using data from Flashbots Protect. halaburda2025impossibility develop a theoretical model of transaction inclusion in blockchains that explicitly incorporates competition among block builders and intermediaries responsible for routing users’ transactions. They show that equilibrium outcomes are centralized, with in one or a few entities controlling transactions inclusion. We complement these theoretical arguments by providing empirical evidence on how these intermediaries affect Ethereum users.
In our companion paper pahari2025exclusiveethereumtransactionsevidence, we use the same data to measure the share of transactions that are exclusive to a builder, meaning they appeared only in blocks proposed by the winning builder during the auction cycle. This question is particularly important given the high concentration in the builders’ market: the top builder (Titan) produces about 50% of all blocks, while the top three account for more than 90%. [^7] For up-to-date statistics, see https://www.relayscan.io/. Buildernet is a joint project by Flashbots (who operated an independent builder until mid-2025) and Beaverbuild (who both participate in Buildernet and run their own builder). We find that between 77% and 84% of the total value of transactions included on-chain is attributable to exclusive transactions, confirming that exclusivity is the main driver of builders’ profits and of PBS auction outcomes.
Finally, to the best of our knowledge, yang2025decentralization is the only other paper that examines the content of non-winning blocks. They study builder competition by comparing bids with block revenues (i.e., total fees paid by the transactions included in a proposed block), but do not analyze individual transactions in non-winning blocks.
The remainder of the paper is organized as follows. The next section describes our dataset and provides additional summary statistics. The following section provides an in-depth analysis of the auction cycle leading to the addition of Block 21322649 to the Ethereum blockchain. We then extend our analysis to all auction cycles in our dataset. The last section concludes.
2 Dataset Description and summary statistics
As already mentioned, our primary dataset includes all blocks (winning and non-winning) submitted via a relay called “Agnostic relay” during the study period. [^8] Founded in 2022, Agnostic relay held 20% of the relay market in 2023, declining to around 5% during our study period. We complement our primary dataset with several publicly available data sources:
  • Winning blocks for our study period, including those not supplied by Agnostic relay. Additionally, we track 10 winning blocks beyond our study period to determine whether transactions in our primary dataset were eventually included on-chain.
  • Hashes of blocks submitted to other relays (Flashbots, BloXroute, Manifold, Eden, Ultra Sound, SecureRpc, and Aestus), along with the name of the builders submitting them. Here, a block’s hash serves as its unique identifier, as two blocks that differ in a single transaction have different hashes. We can, therefore, track whether a block in our primary dataset was also submitted to other relays.
  • Data on transactions submitted via the public mempool, obtained from https://mempool.guru/.
  • Data on transactions submitted via two private mempools, MEV Blocker and Flashbots Protect, obtained from https://Dune.com.
  • Second-by-second price data from Binance (opening price).
We also simulate the non-winning blocks in our dataset to assess how the execution of a transaction — particularly a swap on a blockchain-based financial market — varies depending on the block in which it is included. We simulate the block using an Ethereum Archive Node running the Erigon client, forking at the appropriate block heights.
2.1 Blocks
Ethereum subdivides time into 12-second slots, during which one block at most can be added to the blockchain. For our purposes, each slot corresponds to a bidding cycle. Slot numbers differ from block numbers; for example, our dataset covers all blocks submitted via the Agnostic relay from slot 10,534,387 to 10,534,425, corresponding to winning blocks 21,322,622 to 21,322,660. For ease of exposition, in what follows, we use winning block numbers whenever possible. Also relevant is that a relay will not hold an auction for a given slot if the proposer assigned to that slot is not registered with the relay. This situation appears in our data: out of the 39 slots we consider, we observe no bids for 11 bidding cycles because the proposers were not connected with the Agnostic relay. [^9] Specifically, we observe no bids for the auction cycles that led to the inclusion of blocks 21,322,624; 21,322,625; 21,322,627; 21,322,629; 21,322,633; 21,322,634; 21,322,636; 21,322,644; 21,322,646; 21,322,647; 21,322,651. For the remaining bidding cycles for which we have data, we find that the blocks submitted to Agnostic relay constitute 28.3% of all blocks submitted to all major relays.
Each block in our dataset has two timestamps: received at (the time the relay received the block from the builder) and made available at (the time the relay made the block available to the proposer). The difference between these timestamps arises because relays simulate blocks to verify their validity before making them available to the proposer. [^10] Some blocks are treated “optimistically” and made available to the proposer immediately, with simulation performed in the background. If the simulation later fails, the builder’s next submission will not be treated optimistically. In our dataset, such “optimistic” blocks constitute only 3.8% of the sample. The simulation delay is non-negligible: the median time between the two timestamps is 0.76 seconds, with the 75th percentile at 1.5 seconds. In what follows, we focus on the received at timestamp, as it more accurately reflects the key metrics we study, such as when a builder submits a new block or when a transaction first appears in a block.
As previously noted, our primary dataset includes 15,097 blocks submitted across 28 bidding cycles. The number of blocks per slot ranges from a minimum of 220 to a maximum of 951, with a mean of 539.18 and a median of 509. We are able to match 14,043 of these blocks to 23 known builders. The most active builders in our dataset are Titan Builder (7,024 blocks, 46.5%), Rsync Builder (2,259 blocks, 15.6%), Flashbots (1,936 blocks, 12.8%), and Beaverbuild (1,418 blocks, 9.4%). There is a substantial overlap between blocks submitted to Agnostic and those submitted to other relays: 12,895 blocks in our primary dataset were also shared with at least one other relay.
However, when examining the proportion of blocks submitted by the top builders across all relays, we note an imbalance in our primary dataset. Across all relays, Titan Builder, submitted 13,679 blocks (25.6%), Rsync Builder 7,936 blocks (14.9%), Flashbots 2,626 blocks (4.9%), and Beaverbuild 21,447 blocks (40.2%). Hence, Beaverbuild submits proportionally fewer blocks to Agnostic Relay than to other relays. We acknowledge this imbalance as a limitation of our data.
To explore why builders submit fewer blocks to Agnostic Relay than to other relays, we examine the timestamps associated with the blocks that were shared between Agnostic and at least one other relay. We find that only 5% were received by Agnostic before any other relay. More specifically, among the blocks received by both Agnostic and Ultrasound, just 1% of blocks were received by Agnostic first; between blocks received by both Agnostic and BloXroute only 15% were received by Agnostic first. Agnostic, therefore, has a latency disadvantage relative to the other major relays, which may explain why it receives fewer blocks from some builders.
2.2 Transactions
The first observation is that the set of proposed transactions is very different from the subset of these transactions that appear in winning blocks. The vast majority (87.7%) of the transactions in our primary dataset that appear in winning blocks are users’ transactions, and only a few are searchers’ transactions. [^11] We classify an address as a searcher if it trades on decentralized exchanges (DEXes) using a smart contract whose source code is not disclosed on https://Etherscan.com. In other words, part of the searcher’s logic is embedded in a smart contract but hidden by making only its bytecode available (the bytecode is the compiled, machine-readable version of the smart contract). Swapping on a DEX is necessary for all common types of “searching,” such as sandwich attacks, arbitrage, backrunning, and triggering liquidations on lending protocols. Also, only few (14.4%) of transactions included in winning blocks are swaps. Instead, among the set of proposed transactions, about half are swaps. Also, about half of proposed transactions are from searchers.
To be included in a block, a transaction should include either a fee or a payment to the “Fee Recipient” address, a block-level variable set by the builder (typically an Ethereum address they control). The sum of the transfers to the “Fee Recipient” address the fees attached to the transactions in a block is the block’s potential revenues, that is, the paiment that a builder receives if the block is successfully added to the blockchain.
Just as searchers and users pay builders to include their transactions, builders pay proposers by appending a fee reception transaction at the end of each block. In our primary dataset of 15,097 unique blocks, 13,848 (91.7%) contain a fee reception transaction, totaling 12,205 unique fee reception transactions (some blocks differ in content but include the same fee reception transaction). The remaining blocks — those without a fee reception transaction — are low-value blocks in which the builder sets the fee recipient variable directly to the proposer’s address. In these cases, the bid is the sum of payments sent to the fee recipient address. In what follows, we analyze builder bidding behavior separately from users’ and searchers’ activity. To do so, we label “fee reception transactions” as “bids” so that when we refer to “transactions,” we mean transactions that are not fee reception transactions. This leaves us with 10,793 unique transactions excluding bids. Of these, only 5,576 were eventually included in a winning block—either in the cycle in which the transaction first appeared or in a later one.
A notable feature of our primary dataset is that 2,981 transactions — 35% of all unique transactions — originate from just two addresses. The first address1212120x68d3A973E7272EB388022a5C6518d9b2a2e66fBf. is responsible for 1,873 transactions, of which only 31 were included in a winning block. Notably, only 6 (out of 1,873) transactions from this address appear in blocks not built by Titan, and all 6 are simple token transfers to Binance. The second address1313130x51C72848c68a965f66FA7a88855F9f7784502a7F submitted 1,108 transactions, with just 97 included in a winning block. Only 249 (out of 1,108) transactions appear in blocks not built by Rsync, and only 4 appear in blocks not built by either Rsync or Titan. These 4 transactions are also simple token transfers, including two to addresses associated with Wintermute, an algorithmic trading fund that operates Rsync Builder. Based on these patterns, we refer to them as “Titan-bot” and “Rsync-bot”. [^14] Using a different methodology, heimbach2024non also identified Rsync-bot as a non-atomic arbitrage bot, while Titan-bot did not yet exist during the period covered by their study. wu2025measuring also identifies Rsync-bot and Titan-bot as non-atomic arbitrage bots “owned by” or “exclusive to” their respective builders. Almost all transactions from these bots involve token swaps on decentralized exchanges (DEXs): 1,864 for Titan-bot and 1,095 for Rsync-bot. We later show that these bots frequently trade on the same DEX pools within the same auction cycle. Moreover, when they trade the same token pairs in the same direction, their swaps are almost always mutually exclusive, in the sense that no proposed block contains transactions from both bots. [^15] There is a single exception to this pattern: a Titan-bot transaction selling approximately USD 700 of MOG Coin for ETH appears in some blocks alongside an Rsync-bot transaction selling approximately USD 7,000 of MOG Coin for ETH. Taken together, this trading behavior indicates that both bots primarily engage in non-atomic arbitrage—that is, arbitrage between on-chain and off-chain markets—and frequently compete for the same opportunities.
To further support our identification of “Rsync-bot” and “Titan-bot,” we analyze whether these searchers submit different transactions to different builders. Using transaction hashes, we find no case in which Rsync-bot sent the same transaction to both Rsync and Titan. However, when comparing transaction logs — i.e., the sequence of operations executed by the transaction and their outputs — we find 203 such cases. The discrepancy between transaction hashes and logs arises because Rsync-bot modifies the fee offered depending on the builder: on average, it pays 18% more to Rsync than to Titan for executing the same transaction. Moreover, we find that transactions sent exclusively to Rsync often include a provision that renders them non-executable if the block’s fee recipient is not Rsync. This behavior further supports our interpretation that Rsync-bot is integrated with Rsync and that the 18% fee differential represents the bot’s targeted profit margin when interacting with non-integrated builders. To perform the same analysis for Titan-bot, we consider winning blocks for our study period that are not in our primary dataset (because they were not submitted via Agnostic relay). Doing so, we find 20 transactions originating from Titan-bot included in blocks won by Beaverbuild. In 3 cases, the same transactions also appear in our dataset in blocks built by Titan. When comparing transaction logs, we find another two transactions in blocks built by Beaverbuild and Titan. In both cases, Titan-bot paid more for inclusion (by 43% and 23%), when the transaction was sent to Titan than Beaverbuild.
Swaps on decentralized exchanges (DEXes) are also the most common type of transaction across all unique transactions in our primary dataset. We identify 5,282 swap transactions, of which 3,915 occur on Uniswap V3, 1,222 on Uniswap V2, 142 on Sushiswap, and 102 on Pancakeswap. [^16] We observe 99 transactions that interact with multiple DEXes, which is why the total number of swaps on individual DEXes exceeds the number of unique swap transactions. Despite their prevalence, only 808 swap transactions — approximately 15% of all swap transactions — were ultimately included on-chain. A small number of addresses, all belonging to searchers, account for the majority of swaps: the top four addresses are responsible for 3,649 swap transactions (about 70%), with Rsync-bot and Titan-bot alone responsible for 2,955 (56%). Swaps are also concentrated among a few pools: the top 11 DEX pools account for 2,876 transactions (54.4%).
We also identify 794 swap transactions that interact with 41 distinct DEX routers, services that optimize swap execution by splitting trades across multiple liquidity pools and/or using intermediate assets. The most frequently used routers are the Uniswap Universal Router (190 transactions), Uniswap V3 Router (172), 1inch V5 Router (52), and Uniswap V2 Router (42). Of these, 581 transactions were included in a winning block, implying that the majority of swaps included on-chain were routed through a DEX router. The remaining 213 router-based transactions were not included on-chain, and the vast majority of these (151) originate from a single user address, 0xf5213a6a2f0890321712520b8048d9886c1a9900. This user submitted 153 transactions calling the Uniswap V3 Router, of which only two were ultimately included in a winning block.
3 Block 21322649
We now turn to a detailed empirical analysis of a specific auction cycle—namely, the one that resulted in the inclusion of block 21,322,649 on the Ethereum blockchain. This cycle is noteworthy for two reasons. First, bid data across all relays reveals intense competition between Titan Builder and Rsync Builder, with Rsync ultimately winning the auction (see Figures 1). Second, Rsync-bot and Titan-bot actively compete in the same DEX pools, and account for 94%–97% of the total revenue generated by blocks in which they are included. We can therefore closely examine the dynamics of the bidding cycle by analyzing the two bots’ behavior and their respective builders’ bidding strategies.
Figure 1: Builders’ bids across all relays during the auction cycle for Block 21322649
3.1 Description of the data for the auction cycle of block 21,322,649
For this auction cycle, our primary dataset contains 669 blocks and 504 unique transactions, of which 299 (59%) are swaps. Only 21 swaps (8 from users and 13 from searchers) were included in winning block 21,322,649, which was selected approximately 13 seconds after the previous winning block. An additional 8 swaps were included in winning block 21,322,650, 3 more in winning block 21,322,651 and 1 more in winning block 21,322,652. Notably, all the swaps included in these later blocks are from users. A total of 181 transactions are from Titan-bot and 42 are from Rsync-bot, all of which are DEX swaps. None of the Titan-bot transactions were included in a winning block. In contrast, 5 transactions from Rsync-bot were included, all in the winning block 21,322,649.
Figure 2: Number of Unique transactions, blocks and swaps vs ETH price change on Binance, during the auction cycle for Block 21322649
(a) Swaps traded volume per block
(b) Percentage of the total DEX volume attributable to Rsync-bot and Titan-bot
(c) Number of Unique Public, Private, and Exclusive transactions
(d) Fees earned from different types of transactions (in ETH)
Figure 3:
As the bidding cycle progresses, we observe a sharp increase in the number of swaps (see Figure 2). This is accompanied by a rise in volume swapped in each block (see Figure 3(a)), as well as in the share of swap volume attributable to Titan-bot and Rsync-bot (see Figure 3(b)). The figure also shows the rate or arrival of different types of transaction, by distinguishing between public transactions (i.e., those that appear in the public mempool), private transactions (i.e., those that do not appear in the public mempool but appear in blocks by multiple builders) and exclusive transactions (i.e., those that do not appear in the public mempool and appear in blocks by a single builder). [^17] Throughout the paper, in the case of public transactions, the relevant timestamp is the first time in which a transaction appears in a proposed block, not the time at which it appears in the public mempool. In contrast to public transactions, which arrive at a relatively uniform rate throughout the bidding cycle, private and, especially, exclusive transactions tend to arrive toward the end of the cycle (see Figure 3(c)). Despite arriving later, exclusive transactions are the most valuable ones, as reflected in the distribution of fees paid by different transaction types (see Figure 3(d)).
(a) Bids by builders
(b) Block revenues by builders
(c) Perc. of block value kept by Titan
(d) Perc. of block value kept by Rsync
Figure 4: Bids and value of blocks for different builders (note the changes in the time scale)
Of the 669 proposed blocks in our primary dataset, 413 were submitted by Titan Builder and 123 by Rsync Builder. These two builders also account for all of the “high-paying” blocks, as shown in Panel (a) of Figure 4. Panel (b) shows that the revenues from the submitted blocks — measured as the sum of payments to the builder — closely track the builders’ bids over time. The remaining panels of Figure 4 examine the portion of block revenues that Titan and Rsync retain versus those paid to the proposer as bids. These panels show that, as the bidding cycle progresses, both builders reduce the share of revenue they retain, eventually approaching zero. The convergence to zero is faster and more pronounced for Rsync (Panel d) than for Titan (see Panel c).
3.2 Disparity in Transaction Execution
We now turn to study whether the execution of a transaction changes depending on which builder includes it in a block. For instance, a token swap on a DEX may yield different outcomes depending on the transactions that precede it. We focus on swaps on major DEX protocols — Uniswap V2, Uniswap V3, Pancakeswap, and Sushiswap — that are included in multiple proposed blocks. We compare their execution across blocks using transaction logs, which capture the sequence of contract calls and resulting outputs. This enables us to detect differences in how the same transaction is executed in different blocks.
We begin by examining swap transactions with identical transaction hashes that appear in multiple blocks. As previously discussed, searchers typically modify either the fee or the swap amount during the bidding cycle, resulting in each of their transactions having a unique hash. Consequently, transactions with the same hash across different blocks are likely to originate from users rather than searchers.
The first dimension along which a transaction’s execution may vary depending on the block in which it is included is its execution speed. We identify 12 user swaps that first appeared during the bidding cycle for block 21,322,649 and were still present in the subsequent cycle. Of these 12 delayed transactions, 5 were exclusive to Titan during the 21,322,649 cycle, meaning that they appeared only in blocks built by Titan. The remaining 7 were present in blocks by Flashbots, Buildernet, and Titan. [^18] These 7 transactions are different from the 3 transactions identified in the previous subsection because these 3 transactions were present only during the bidding cycle for block 21,322,649. In the following cycle, 3 of the 5 transactions initially exclusive to Titan remain exclusive, while the other 2 appear in blocks by multiple builders, having been submitted via both MEV Blocker and Flashbots Protect. All 9 transactions that were not exclusive to Titan in the second cycle are included in the winning block for that cycle (21,322,650), which was built by Beaverbuild. 2 of the remaining 3 transactions are included in block 21,322,651 (built by Titan), for which we do not have data other than the winning block. [^19] The hashes of those transactions are
0x758f12756ba2a91d417a940e311f4500c449e464e0233a4b896203228302af22
0x100825b825744accd39b532d2f7c98c5d8472154b21092294e83f227e9d6d2ff
0x74b6b90aa9ec32adc9e35c15fbd1cf5af37a8ee4d43141785ed21f3bef3f6d8b
The last transaction is a user’s swap that remains exclusive to Titan and is finally included in winning block 21,322,652 (also built by Titan). [^20] Its hash is 0x74b6b90aa9ec32adc9e35c15fbd1cf5af37a8ee4d43141785ed21f3bef3f6d8b Hence which builder has access to which transactions affects not only whether transactions are included on-chain, but also when they are included.
The second dimension along which a transaction’s execution may vary depending on the block in which it is included is whether the transaction succeeds or fails. For the auction cycle under consideration, we identify seven transactions that fail in at least one block: six of these fail in all blocks in which they appear and four of those always fail also when we simulate them at the top of each block in which they were included, suggesting that they were erroneously constructed or misconfigured. At least for this bidding cycle, whether a swap fails does not seem to depend on the block in which it was included.
Finally, the same swap transaction may be executed at different prices depending on the block in which it is included. Among the 12 users’ swaps observed across multiple bidding cycles, only 3 exhibit variation in execution price. Surprisingly, for those 3 swaps, execution quality is better in the later cycles. We also identify 7 users’ transactions with different execution prices across blocks within the bidding cycle for block 21,322,649. Two swap ETH for MogCoin and receive the best execution in a block built by Rsync, where they are preceded by a swap by Rsync-bot selling MogCoin. However, we also find 3 transactions selling MogCoin that achieve worse execution in blocks built by Rsync and Titan (compared to other builders). Overall, transactions that trade in the opposite direction as the two bots tend to receive better execution in blocks built by Rsync and Titan, while those swapping in the same direction as the two bots receive worse execution in blocks built by Rsync and Titan.
3.3 Competition between Rsync-bot and Titan-bot.
Having established that the identity of the winning builder affects the quality of users’ transaction execution, we now turn to the behavior of Rsync-bot and Titan-bot. As noted earlier, during this auction cycle, these two bots account for between 94% and 97% of the total revenues of submitted blocks, making them one of the main drivers of the auction’s outcome.
We identify 42 unique transactions from Rsync-bot, interacting with five DEX pools: USDC/WETH, USDT/WETH, MOGCOIN/WETH (2 pools), and WBTC/cbBTC. For Titan-bot, we observe 181 unique transactions across five DEX pools: USDC/WETH, USDT/WETH, MOGCOIN/WETH (2 pools), and MATIC/WETH. Notably, both Rsync-bot and Titan-bot execute swaps that buy WETH for MOGCoin and buy stablecoins (USDC and USDT) for WETH. This suggests that the two bots are competing for the same arbitrage opportunities between decentralized exchanges (DEXes) and off-chain centralized exchanges.
Our data allow us to study the competition between Rsync-bot and Titan-bot by comparing:
  1. 1.
    the trading volume of each bot and how it evolves over the course of the auction cycle;
  2. 2.
    the fees each bot pays and how they change during the auction cycle;
  3. 3.
    the price paid on the DEX, both in terms of the raw execution price (based on token in/out amounts) and the effective price net of the inclusion fee.
Furthermore, by interpreting the fee attached to each transaction as the expected profits in case that transaction is added on chain, we can infer a “risk-adjusted implied price” on the centralized exchange (CEX) side of the arbitrage path—a component that is typically unobservable. This method provides new insights into the off-chain leg of arbitrage activity.
More precisely, suppose a bot buys a given token for ETH on a DEX at price $p_{\text{DEX}}$ (quoted in ETH) and sells the same token for ETH on a centralized exchange (CEX) at price $p_{\text{CEX}}$. The bot’s profit from the arbitrage trade, in ETH, is:
$\pi=v\cdot(p_{\text{CEX}}-p_{\text{DEX}}),$
where $v$ is the volume of the trade, expressed in units of the non-ETH token. While we do not observe $p_{\text{CEX}}$ directly, we do observe the fee paid for block inclusion, $f$, which we interpret as the bot’s risk-adjusted expected profit. Substituting this into the equation above allows us to compute the implied risk-adjusted CEX price:
$p_{\text{implied}}=p_{\text{DEX}}+\frac{f}{v}.$
If, instead, the trade is the opposite direction—selling a token for ETH on the DEX and buying it on the CEX—the formula becomes:
$p_{\text{implied}}=p_{\text{DEX}}-\frac{f}{v}.$
Note that all prices here are quoted in ETH. In some cases, such as trades involving stablecoins, we may prefer to express prices in terms of the other token. This simply requires inverting the expressions above. Finally, each observation of $p_{\text{implied}}$ has a timestamp equal to the timestamp of the block including the swap transaction representing the DEX leg of the trade.
(a) WETH bought on the DEX
(b) Fees paid for inclusion (in ETH)
(c) MogCoin / ETH implied DEX price
Figure 5: Competition between Rsync-bot and Titan-bot on the WETH/MOG Uniswap v3 pool.
(a) Amount of ETH sold on the DEX
(b) Amount of ETH sold on the DEX
(c) Fees paid for inclusion (in ETH)
(d) Fees paid for inclusion (in ETH)
(e) USDT / ETH implied DEX price with Binance price
(f) USDC / ETH implied DEX price with Binance price
Figure 6: Competition between Rsync-bot and Titan-bot on the WETH/USDT and WETH/USDC Uniswap v3 pools.
Figure 5 presents the results of our analysis for MOG/WETH swaps on the Uniswap V3 pool, and Figure 6 replicates the analysis for swaps on the USDT/WETH and USDC/WETH Uniswap V3 pools. The first observation is that, despite the fact that the amount swapped is similar on the three markets, the fee paid by the bots for swapping on the MOG/WETH pool is orders of magnitude larger than that paid for swapping in the other two pools. The MOG/WETH arbitrage opportunity is therefore much more valuable than the other two and drives the value of winning this specific auction cycle. Focusing on the MOG/WETH swaps, we note that Titan-bot maintains a constant trade volume throughout the auction cycle, but its fees increase. We interpret this as evidence that the risk associated with the arbitrage opportunity decreases as the auction progresses, thereby increasing its expected profit. Rsync-bot instead increases both the volume traded on the DEX and its fees as the auction cycle progresses.
Finally, in both Figure 5 and 6, we adjust the base currency in the graphs for the implied centralized exchange (CEX) price so that higher prices are better for the bots. The main result here is that, for the MOG/WETH arbitrage, Titan-bot’s implied CEX price is 1.9% better than that of Rsync-bot (Note that MOG Coin was not listed on Binance nor Coibase during our study period). For USDT/WETH and USDC/WETH, instead, Rsync-bot seems to have an advantage of approximately 0.17% and 0.02%, respectively. Note also that the implied CEX prices are sometimes above or below Binance’s. Remember that the implied CEX prices are risk-adjusted, that is, the arbitrage risk is reflected in a less favorable implied CEX price. Also, the reported Binance price does not account for trading fees or price impact. These reasons may explain why the bots’ implied CEX prices may be below those on Binance. At the same time, the bots may trade on multiple centralized markets or have their own inventory, which they may use to achieve an even better execution price than Binance’s.
To summarize, we showed that the MOG/WETH arbitrage is significantly more valuable than the other two, and hence, Titan-bot can generate considerably more profits than its competitor. Despite this, Rsync Builder wins the block because, as we already discussed, it increases its bid more aggressively than Titan.
4 Entire Dataset
We now extend our analysis to all bidding cycles in the dataset. We first examine differences in transaction execution across blocks and then analyze the behavior of Titan-bot and Rsync-bot.
4.1 Time of inclusion in a winning block
We identify 1,288 transactions that were present in multiple bidding cycles. These transactions appeared in a bidding cycle before the winning block was selected, were not included in the winning block, and appeared again in the subsequent bidding cycle. Among these, 871 were present for two cycles, 198 for three cycles, and 108 for four cycles before being included in a winning block (see Figure 7). Remarkably, one transaction remained present for 2,437 bidding cycles before inclusion; it was exclusive to the builder payload.de and was ultimately included only when payload.de won. In addition, 28 transactions appeared in multiple cycles but were never included in a winning block. Among the transactions present in multiple bidding cycles, only 17 are searchers’ transactions. Of the remaining transactions, 701 first appeared in our dataset as public (i.e., present also in the public mempool), 189 as private (i.e., present in blocks by multiple builders but not in the public mempool), and 381 as exclusive to a single builder (with 374 exclusive to Titan Builder).
Figure 7: Empirical CDF of the waiting time to on-chain inclusion, measured in blocks, for users’ transactions.
Hence, the vast majority of searchers’ transactions appear in only a single bidding cycle and are either included in the winning block or discarded. Transactions present across multiple bidding cycles are predominantly users’ transactions. Given that our dataset includes 5,968 user transactions, we estimate that approximately 21% users’ transactions are delayed. This delay is partly explained by the fact that 30% of these transactions are, at least initially, exclusive to a single builder who does not win.
In addition, 12 transactions that were delayed because exclusive eventually appear in the public mempool. Nine of these were exclusive for two cycles and three for three cycles. In four cases, after appearing in the public mempool, these transactions were added on chain by the same builder that had them as exclusive to start with (Titan). All appear to be from users: three used the 0x router and four the 1inch router. The existence of these transactions is problematic, as they were first delayed and later exposed in the public mempool, leaving users vulnerable to attacks (and one of them appears to have been sandwiched). [^21] These transactions are: 0x1daea1584d684385fb25209bad3a49c54a1e2c500e27543334c7b3e695a12ffb, 0x6605c1c36fcd3d696844519d6ff7199fe4ddd34fb4f3e0ed4da71aedfcf1c633, 0x3dadbab8e85b6e48a2c33f9029ea4ebebdd3def8b9d907faff15ef81527de968, 0x9c7364eddb94fe659bc8104585ff86a8a234480fe00a9f9aa3eb410e572ebdf2, 0x9d476dbc524651cbe26bcafab99f6fe40a823a5588fa5cd5002e5b414a8a2f8c, 0x417440b083bdcb529663c6615c32d89a92beb7863f1a5d6b12bca5b94f4a0f87 (sandwiched according to https://www.zeromev.org/block?num=21322643), 0xe56caee27e79dee8a4c7ed6f945ccbc2cc733d761e19bd68329dc364c61e73a7, 0xae5b88bc0bffa04811eda0dfdcd7224599e2757339e18b2d57f4f40d61902a4c, 0xd41139b67a193045b4d60266c084f6dadd0cb95dccd7b2a06ea68653d6a313ec, 0x597fc0cb0610d1e3ea2baf0f6f90d0d741a6429bfdd9b375bfd2493b70d41dcd, 0x44e8105d5c672195877dc1ed970511e0f0a15fc25cbcafe71d51d428d54c215d, 0xd74718e6d8d58bc33334c0c2f6642d9dec49496329329d8fdcb1e698e81abc94.
Unfortunately, our data do not allow us to determine with certainty which wallet was used to create a given transaction, which private mempool (or other service) was used to relay it to builders, or how the transaction was propagated through the public mempool. This information is simply not recorded in Ethereum blocks. Nonetheless, a plausible explanation is that some private mempool operators share transactions exclusively with a single builder, reverting to the public mempool if that builder fails to win two or three consecutive auctions.
4.2 Successful execution
A transaction may execute successfully or fail depending on the transactions that precede it. Because a given transaction may be preceded by different sets of transactions across the blocks in which it appears, the same transaction may execute successfully in some proposed blocks but fail in others.
To study this possibility, we consider each transaction–block pair and classify it as a failure if its simulated execution yields status = 0, or if it produces no transaction log despite generating one when included in another block. A success is defined as any transaction–block pair not classified as a failure. Under this definition, failures include both outright execution failures and reverts, such as a swap reverting because its execution price would fall below the user’s specified minimum acceptable price. By contrast, a success indicates that the transaction would have executed as intended had the corresponding block been added to the chain.
We identify 149 transactions that always fail and 49 transactions that succeed in some blocks but fail in others. Of these 49 transactions, 44 are swaps: 10 generated by searchers and 34 by users. Because our dataset contains 10,793 transactions—of which 4,565 are searchers’ swaps and 717 are users’ swaps—this implies that virtually all searcher swaps always execute successfully, while 3.3% of users’ swaps fail in at least one of the blocks in which they appear. Virtually all remaining transactions either always succeed or always fail. [^22] A subtlety arises because transactions that always fail does not generate any log and cannot be classified by type. As a result, identifying a transaction as a swap requires that it execute successfully in at least one block.
We therefore focus on user swaps that sometimes fail and plot the empirical CDF of their fraction of failures (Figure 8). Approximately 50% of these transactions fail in at least half of the candidate blocks in which they appear, and roughly 30% fail in more than 80% of such blocks. The main takeaway is that the probability a user’s swap fails is highly skewed: the vast majority of users’ swaps (96.7%) never fail, while a small fraction fail very frequently—1.7% fail in at least half of the blocks in which they appear, and 1.1% fail in more than 80% of such blocks. This skewness indicates that execution risk is concentrated among a small subset of user swaps, rather than being evenly distributed.
Figure 8: Empirical CDF of the fraction of candidate blocks in which a given user swap fails, conditional on the swap sometimes failing and sometimes succeeding.
We then perform a regression analysis restricted to swaps that both fail and succeed, which account for 318,742 transaction–block observations. For each swap $i$ appearing in block $k$, we construct a binary outcome variable equal to one if transaction $i$ executes successfully in block $k$, and zero otherwise. We then regress this outcome on the following explanatory variables:
  • Time since Last Block: the time elapsed since the previous block, measured at the moment the block was submitted for inclusion.
  • Tx index: the position of the transaction within the block, where lower values indicate earlier placement (and thus earlier execution). For example, Tx index = 1 indicates the first transaction in the block, Tx index = 2 the second, and so on. [^23] We also tested a normalized specification in which transaction position is scaled from 1 to 100, with 1 indicating the first and 100 the last transaction in the block. All results are robust to this alternative.
  • A dummy variable indicating whether the block was built by Titan.
  • A dummy variable indicating whether the block was built by Rsync.
  • A dummy variable indicating whether the block contains a transaction from Titan-bot.
  • A dummy variable indicating whether the block contains a transaction from Rsync-bot.
We also include transaction fixed effects: that is, a dummy variable for each unique transaction hash. These control for transaction-specific characteristics and capture the average likelihood that a transaction executes successfully across all blocks in which it appears. The coefficients on the other variables then measure how the probability of successful execution varies depending on block timing, builder identity, transaction ordering, and the presence of the two bots. Finally, we construct two subsamples: swaps in the same direction as those by Titan-bot and Rsync-bot, and swaps in the opposite direction. To do so, for each swap by either Titan-bot or Rsync-bot, we record the target pool, its direction, and the auction cycle (noting that, in our data, all bots’ swaps on a given pool and during the same cycle are in the same direction). We then identify all other swaps occurring in those same pools and cycles, and divide them between those “in the same direction” as the bots and those “in the opposite direction” as the bots.
Table 1 reports three regressions, each on a different samples. The first regression yields mixed results: being included in a block built by Titan or Rsync reduces the probability of successful execution, but the presence of transactions from the corresponding bot increases it. In contrast, the second and third regressions reveal a clearer pattern. Swaps in the same direction as the bots are significantly more likely to fail when they are included in blocks built by Rsync or Titan, especially if those blocks also contain transactions from the bots themselves. The opposite holds for swaps in the opposite direction: their likelihood of success increases in blocks which also contain transactions from the bots. We also observe that swaps included later in a block (i.e., with a higher transaction index) are slightly less likely to fail. This result is, however, difficult to interpret, as transaction ordering within a block is determined by the builder. Finally, transactions are more likely to fail when included in blocks submitted later in the bidding cycle.
Using the estimated coefficients, we perform back-of-the-envelope calculations to quantify the effects. Compared to inclusion in a block built by a non-searcher-integrated builder, a swap in the same direction as Titan-bot or Rsync-bot is approximately 18% less likely to execute successfully in blocks that also contain a transaction from these bots (remember that almost all blocks with transactions from these bots are built by the corresponding builder). By contrast, a swap in the opposite direction is approximately 1% more likely to execute successfully under the same conditions.
Table 1: Probability of Success
All Swaps Same Dir. as Bots Opp. Dir. as Bots
Time Since Last Block -0.0038*** -0.0217*** -0.0012***
(0.0001) (0.001) (0.000)
Tx Index 0.0001*** 0.0006*** 5.74e-06
(0.0000) (3.18e-05) (9.23e-06)
Is Titan Builder -0.0078*** -0.1341*** -0.0144***
(0.0008) (0.005) (0.001)
Is Rsync Builder -0.0148*** -0.1592*** -0.0021
(0.0006) (0.006) (0.002)
Has Titan-bot tx 0.0024*** -0.0423*** 0.0286***
(0.0005) (0.004) (0.001)
Has Rsync-bot tx 0.0075*** -0.0301*** 0.0095***
(0.0005) (0.004) (0.001)
Observations 318742 17224 26705
R-squared 0.744 0.556 0.493
Tx fixed effect yes yes yes
    $p<0.1$,
    $p<0.05$,
    $p<0.01$
Note: Standard errors in parentheses. Column titles refer to subsample type.
4.3 Execution price
The way a swap is included in a block may also influence its execution price. To investigate this possibility, we examine all swap in blocks where they executed successfully and the execution price can be computed. This yields 272 unique transaction hashes, 112,747 transaction-block observations, and 342,528 swap-block observations, which will be our unit of analysis (note that many transactions perform multiple swaps).
For each swap $i$, we compute its average execution price across all blocks, denoted by $p_{\text{avg},i}$. We then calculate, for each transaction-block observation, the percentage deviation from this average:
$p_{dev,i,k}\equiv\frac{p_{i,k}-p_{\text{avg},i}}{p_{\text{avg},i}}\times 100$
where $p_{i,k}$ is the execution price of swap $i$ in proposed block $k$. We adjust the base currency so that higher values of $p_{dev,i,k}$ always reflect better execution from the user’s perspective. The distribution of $p_{dev,i,k}$ across all $i$ and $k$ has a median of $1.41\times 10^{-14}$, a minimum of 10.53-10.53 and a maximum of $13.47$ basis points.
Figure 9: The CDF of the standard deviation of the $p_{dev,i,k}$.
For each transaction $i$, we compute the standard deviation of $p_{\text{dev},i,k}$. This yields one observation per swap, capturing the dispersion of its execution price across candidate blocks. Figure 9 plots the distribution of these standard deviations. As before, the distribution is extremely skewed: the median is 0.043, while the mean is 0.27. Moreover, for 15% of swaps the standard deviation of $p_{\text{dev},i,k}$ is below $10^{-9}$, whereas for 7% of swaps it exceeds $1$. For the sake of interpreting these magnitudes, suppose that for each swap $i$ the distribution of $p_{\text{dev},i,k}$ is normal. Under this assumption, for half of the swaps in our sample there is less than a one-third probability that the execution price deviates by more than 4.3 basis points from $p_{\text{avg},i}$, while for 7% of swaps there is more than a one-third probability that the execution price deviates by at least 100 basis points from $p_{\text{avg},i}$.
Also here, we resort to a regression analysis to decompose such variation. Table 2 reports the coefficients from three regressions of $p_{dev,i,k}$ on the same set of explanatory variables and subsamples used in the success/failure analysis. As with the previous table, the regression on the full sample yields ambiguous results: being included in a block built by Titan is associated with slightly worse execution prices, while the presence of a transaction from Titan-bot in the same block improves execution. For Rsync, the pattern is reversed. These opposing effects largely offset each other. In contrast, the regressions restricted to swaps in the same direction as either bot show a clearer trend: execution prices are between 9 and 12 basis points worse when such swaps are included in blocks built by Rsync or Titan (the presence of the two bots has a negligible effect). For swaps in the opposite direction from the bots, results align with those from the success/failure regressions: being included in a block built by Rsync or Titan improves the execution price by between 9 and 27 basis points. The effect remains positive (but smaller) if the block also contains a Rsync-bot or Titan-bot transaction. Additionally, being placed later in the block (i.e., with a higher transaction index) has a small but consistently positive effect on price. Finally, while many coefficients are statistically significant, the regressions explain only a modest share of the variation in $p_{dev,i,k}$, as reflected in the low $R^{2}$ values.
Table 2: Price deviations $p_{dev,i,k}$
All Swaps Same Dir. as Bots Opp. Dir. as Bots
Time since Last Block -0.0022*** -0.0430*** 0.0163***
(0.0005) (0.002) (0.002)
Tx Index 0.0003*** 0.0001*** 0.0007***
(0.0000) (4.01e-05) (4.28e-05)
Is titan builder -0.0053** -0.1206*** 0.0923***
(0.0024) (0.006) (0.007)
Is rsync builder 0.0198*** -0.0924*** 0.2758***
(0.0031) (0.008) (0.009)
Has titan-bot tx 0.0048** 0.0068* -0.0640***
(0.0019) (0.005) (0.006)
Has rsync-bot tx -0.0235*** -0.0038 -0.0402***
(0.0019) (0.005) (0.006)
Observations 342528 15404 25718
R-squared 0.002 0.142 0.073
Tx fixed effect yes yes yes
    $p<0.1$,
    $p<0.05$,
    $p<0.01$
Note: Standard errors in parentheses. Column headers refer to different subsamples.
4.4 Competition between Searchers
In the previous section, we highlighted the importance of competition between Rsync-bot and Titan-bot. It turns out that these two searchers compete intensely throughout our entire study period, and their competition is a key determinant of which builder wins the block: every auction cycle for which we have data includes at least one transaction from these bots, which together account for approximately 60% of the builders’ revenues in the blocks where they appear, underscoring their central role in shaping auction outcomes.
We identify 57 instances in which Titan-bot and Rsync-bot compete for execution in the same DEX pool during the same auction cycle. In 52 cases, one of the tokens swapped is WETH, allowing us to apply the same methodology introduced in the previous section — namely, deriving an implied risk-adjusted CEX price for each bot and different moments during a bidding cycle. To enable direct comparison with observed market data, we further restrict the sample to tokens traded on Binance. This yields 29 auction cycle/DEX pool combinations for which we can compare the bots’ implied prices to contemporaneous Binance prices.
We plot the implied DEX price alongside the Binance price for all 29 cases in Appendix A (note that Binance reports prices at one-second intervals, we use the opening price). As expected, the implied DEX prices generally track the Binance price closely. This correlation is particularly strong in cases involving the USDT/WETH and USDC/WETH pairs, which account for the majority of the sample. A weaker but still visible relationship is observed for the UNI/WETH and AAVE/WETH pairs. In contrast, for the DAI/WETH pair, the implied DEX and Binance prices do not exhibit a consistent relationship, likely because Binance is not the primary centralized exchange for DAI trading. This is confirmed by the fact that the total volume on the DAI/ETH market on Binance during our study period was only 2.5 ETH. [^24] For comparison, total volume on Binance during the study period for the other markets was: 1,487 ETH for the USDC/ETH market, 16,209 ETH for the USDT/ETH market, 1,374 ETH for the WBTC/ETH market, 6.9 ETH for the UNI/ETH market, 36 ETH for the AAVE/ETH Binance market.
We complement the visual analysis with a regression approach. For each second $s$, we compute the average implied CEX price and trade volume for Rsync-bot and Titan-bot by considering all trades executed by each bot between seconds s0.5s-0.5 and $s+0.5$. We denote the average implied CEX price for second $s$ by $p_{\text{implied},s}$. For technical reasons, we restrict our analysis to the USDT/WETH and USDC/WETH trading pairs. [^25] As discussed earlier, Binance prices for some token pairs appear unreliable. More importantly, the number of observations, values of $p_{\text{implied},s}$, and trade volumes vary considerably across tokens. We have 104 observations for WETH/USDC and WETH/USDT combined, but only 24 for WETH/DAI, 18 for WETH/UNI, and 5 or fewer for WBTC/WETH and AAVE/WETH. Moreover, the average value of $p_{\text{implied},s}$ is around 3,500 for stablecoin pairs (USDC, USDT, DAI), but drops to 261 for UNI/WETH, 15 for AAVE/WETH, and 0.037 for WBTC/WETH. Trade volumes also vary: average volume is 320.3 ETH for WETH/WBTC, compared to 43.1 ETH for other assets. As a result, separate regressions on the less common pairs are underpowered, while pooled regressions suffer from multicollinearity due to token-specific heterogeneity. This filtering results in 89 total observations: 43 for Rsync-bot and 46 for Titan-bot. We regress $p_{\text{implied},s}$ on the contemporaneous Binance opening price, as well as its one- and two-second lag and one-second lead. The regression results, shown in Table 3, indicate that only the contemporaneous and one-second lagged Binance prices are significantly correlated with the bots’ implied CEX prices. Among these, the contemporaneous price has the strongest explanatory power. Therefore, the contemporaneous Binance price serves as the most appropriate benchmark for evaluating $p_{\text{implied},s}$.
Table 3: Implied Price on Different Binance Prices
Coefficient Standard Error
Binance Price (0 sec.) 0.79450.7945*** $0.108$
Binance Price (+1 sec.) 0.1491-0.1491 $0.129$
Binance Price (-1 sec.) 0.25130.2513** $0.108$
Binance Price (-2 sec.) $0.0972$ $0.063$
Constant $21.5424$ $23.070$
Observations 89
R-squared 0.998
    $p<0.1$,
    $p<0.05$,
    $p<0.01$
We then proceed to compare each observation of $p_{\text{implied},s}$ to the corresponding contemporaneous Binance price, $p_{\text{binance},s}$. To interpret the execution quality from the bot’s perspective, we distinguish between two cases. If a bot sells ETH on-chain, it must be buying ETH on the CEX; in this case, a lower CEX price is better. We therefore define
$p_{\text{diff},s}\equiv p_{\text{binance},s}-p_{\text{implied},s}.$
Conversely, if the bot buys ETH on-chain we define
$p_{\text{diff},s}\equiv p_{\text{implied},s}-p_{\text{binance},s}.$
In both cases, a higher $p_{\text{diff},s}$ indicates better execution for the bot. We then calculate
$p_{\text{Improvement},s}\equiv\frac{p_{\text{diff},s}}{p_{\text{binance},s}}\times 100$
as the percentage price improvement relative to the contemporaneous Binance price. The minimum, maximum, and average values of $p_{\text{Improvement},s}$ are -0.01428, 0.04655 and 0.01285 respectively.
Table 4 reports the coefficients of two separate regressions. Both include a constant, a dummy variable equal to 1 if the observation is from Titan bot, and time (seconds) since the last block. The difference between the two regressions is that the second includes swap volume (in WETH) in piecewise linear form, where each piece corresponds to 50 ETH. The coefficient on the constant represents the average percentage price improvement ($p_{\text{Improvement},s}$) for a zero-volume swap by Rsync-bot, assuming the block was built at time zero in the bidding cycle. The coefficient on Is Titan bot captures the difference in execution quality between Titan-bot and Rsync-bot. The results indicate that the bots achieve better execution than the Binance price, and that Rsync-bot outperforms Titan-bot. Execution quality seems to improve with trade size, although this result needs to be taken with a grain of salt because there are only 2 trades of volume larger than 150. Using the estimated coefficients, we perform back-of-the-envelope calculations. For a swap of volume lower than 100 ETH the implied CEX price for ETH is approximately 2.8 basis points better than the contemporaneous Binance price for both bots.
Finally, recall that our derivations assume the fee paid by the bots for inclusion equals their risk-adjusted profits expected if their transactions are included on-chain. If this assumption is relaxed, the observed fee provides a lower bound on expected profits: bots may offer builders less than their total expected profits in order to retain part of those profits. In that case, our regression estimates should be interpreted as lower bounds, indicating that the two bots trade on the CEX leg of their arbitrage at prices at least 2.8 basis points better than the contemporaneous Binance price.
Table 4: Regression of percentage price improvement over Binance ($p_{\text{Improvement},s}$) with and without piecewise volume
(1) (2)
Time since Last Block -0.009 -0.001
0.001 0.001
Is Titan Bot -0.0069** -0.007**
0.003 0.003
Volume $\times$ (Volume lower than 50) -0.0007
0.004
Volume $\times$ (Volume between 50 and 100) 0.0003
0.004
Volume $\times$ (Volume between 100 and 150) 0.0066
0.004
Volume $\times$ (Volume greater than 150) 0.0231***
0.008
Constant 0.028 0.0287**
0.018 0.013
Observations 89 89
R-squared 0.080 0.195
    $p<0.1$,
    $p<0.05$,
    $p<0.01$
Note: standard errors in parentheses
5 Conclusions
We study the Ethereum blockchain by analyzing blocks that were proposed for inclusion into the canonical chain but were ultimately discarded. To the best of our knowledge, together with the companion paper pahari2025exclusiveethereumtransactionsevidence, ours is the first systematic analysis of non-winning Ethereum blocks. Our findings illuminate key aspects of Ethereum’s block-building process, with implications for ongoing discussions among practitioners and for the academic literature.
While informative, our results should be interpreted in light of certain limitations. As noted, the dataset underrepresents some major builders (notably Beaverbuild) and spans only eight minutes. To address these gaps, we hope that relays, builders, and other participants in the transaction supply chain will make more comprehensive data available in the future, enabling a deeper and more rigorous understanding of how Ethereum is made.
Appendix A Competition between Rsync-bot and Titan-bot
(a) USDC Pool 1
(b) USDC Pool 2
(c) USDT Pool 1
(d) USDT Pool2
Figure 10: Block 21322626
(a) USDC Pool 1
(b) USDT Pool 1
(c) USDT Pool 2
Figure 11: Block 21322630
(a) USDC Pool 1
(b) USDC Pool 2
Figure 12: Block 21322631
(a) USDC Pool 1
(b) USDC Pool 2
(c) USDT Pool 1
Figure 13: Block 21322648
(a) USDC Pool 1
(b) USDT Pool 1
Figure 14: Block 21322649
(a) USDC Pool 1 Block 21322657
(b) USDC Pool 1 Block 21322658
(c) USDT Pool 2 Block 21322659
(d) USDT Pool 3 Block 21322660
Figure 15: Block 21322657 - 21322660
(a) DAI Block 21322635
(b) DAI Block 21322659
(c) DAI Block 21322660
(d) DAI Block 21322660
(e) DAI Block 21322660
Figure 16: DAI all blocks
(a) UNI Block 21322628
(b) UNI Block 21322639
(c) UNI Block 21322643
(d) UNI Block 21322643
(e) UNI Block 21322653
(f) UNI Block 21322653
Figure 17: UNI all blocks. Note that panels (e) and (f) plot the same data but at different time scales.
Figure 18: AAVE Block 21322658
Figure 19: WBTC Block 21322641